### v2

# Coupled Data and Measurement Space Dynamics for Enhanced Diffusion Posterior Sampling

Abstract
Inverse problems, where the goal is to recover an unknown signal from noisy or incomplete measurements, are central to applications in medical imaging, remote sensing, and computational biology. Diffusion models have recently emerged as powerful priors for solving such problems. However, existing methods either rely on projection-based techniques that enforce measurement consistency through heuristic updates, or they approximate the likelihood $p(\boldsymbol{y}\mid\boldsymbol{x})$, often resulting in artifacts and instability under complex or high-noise conditions. To address these limitations, we propose a novel framework called coupled data and measurement space diffusion posterior sampling (C-DPS), which eliminates the need for constraint tuning or likelihood approximation. C-DPS introduces a forward stochastic process in the measurement space $\{\boldsymbol{y}_{t}\}$, evolving in parallel with the data-space diffusion $\{\boldsymbol{x}_{t}\}$, which enables the derivation of a closed-form posterior $p(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t},\boldsymbol{y}_{t-1})$. This coupling allows for accurate and recursive sampling based on a well-defined posterior distribution. Empirical results demonstrate that C-DPS consistently outperforms existing baselines, both qualitatively and quantitatively, across multiple inverse problem benchmarks.
1 Introduction
Inverse problems, where the goal is to recover an unknown signal $\boldsymbol{x}_{0}$ from noisy or incomplete measurements $\boldsymbol{y}$, arise in a wide range of applications, including medical imaging chung2023solving; chung2022score, remote sensing twomey2019introduction; meng2024conditional, and audio signal processing moliner2023solving; saito2023unsupervised. Mathematically, these problems are often modeled as $\boldsymbol{y}=\mathbb{A}\boldsymbol{x}_{0}+\mathbf{n}$, where $\mathbb{A}$ is a known forward operator and $\mathbf{n}$ represents measurement noise. Inverse problems are inherently ill-posed, meaning that, without additional constraints, infinitely many solutions may satisfy the given measurements $\boldsymbol{y}$. As such, incorporating prior knowledge about the underlying signal and the noise model is essential for reliable reconstruction.
One principled approach to addressing this uncertainty is to treat inverse problems in a Bayesian framework, where the goal becomes sampling from the posterior distribution $p(\boldsymbol{x}_{0}|\boldsymbol{y})$. However, accurate and efficient posterior sampling remains a central challenge, particularly when the forward operator $\mathbb{A}$ is ill-conditioned or the measurement noise is significant DPS. Traditional sampling-based methods, such as Markov chain Monte Carlo (MCMC), often struggle with high-dimensional spaces or require careful tuning of proposal distributions song2023pseudoinverse. More recently, diffusion-based techniques have emerged as powerful generative models for high-dimensional data, and several works have adapted diffusion processes to inverse problems by combining a learned data prior $p(\boldsymbol{x}_{0})$ with a measurement-consistency term DPS; song2023pseudoinverse; NEURIPS2024_54fa8255; choi2021ilvr; ye2024decomposed; MCG; DSG; kawar2022denoising; hamidi2025conditional; boys2023tweedie; NEURIPS2024_2f46ef57; NEURIPS2024_03261886; hamididiffusion.
Nevertheless, in standard diffusion-based frameworks, the measurement information must often be retro-fitted into the prior $p(\boldsymbol{x}_{0})$ via projection-based techniques song2023pseudoinverse; NEURIPS2024_54fa8255; choi2021ilvr; ye2024decomposed; MCG; DSG; kawar2022denoising; hamidi2025conditional or approximation of the likelihood $p(\boldsymbol{y}|\boldsymbol{x})$ DPS; boys2023tweedie; NEURIPS2024_2f46ef57; NEURIPS2024_03261886; hamididiffusion (see Appendix˜A for details). Since these methods modify the iterative processes in data and measurement space in an uncoordinated manner, there is no guarantee that the generated samples lie on a valid data manifold while also being consistent with the observed measurements. This often leads to suboptimal reconstructions: the generated samples may either exhibit visual artifacts due to drifting off the data manifold, or fail to resemble the original measurements due to poor measurement fidelity. Examples of both failure modes are illustrated in Figure˜1.
Figure 1: Visualization of reconstructed samples for some linear inverse problems using our proposed C-DPS method, compared with two leading baselines: DPS DPS and Resample resample. C-DPS exhibits less visual artifacts and improved measurement consistency.
To address the limitations of conventional approaches, we propose C-DPS (coupled data and measurement space diffusion posterior sampling), a novel framework that extends the standard diffusion model by introducing a second, parallel diffusion process in the measurement space, alongside the conventional diffusion in the data space. While the data space process follows the traditional forward diffusion from an unknown signal $\boldsymbol{x}_{0}$, the measurement space process begins with the known observation $\boldsymbol{y}_{0}$ and progressively injects noise. This symmetric treatment mirrors the dynamics applied to $\boldsymbol{x}_{0}$, tightly coupling the two spaces throughout the diffusion trajectory. This coupling enables the derivation of a closed-form posterior expression for $p(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t},\boldsymbol{y}_{t-1})$, thereby removing the need for ad-hoc measurement terms or learned likelihood approximations commonly used in prior work.
From a Bayesian perspective, the joint diffusion over both $\boldsymbol{x}$ and $\boldsymbol{y}$ defines a coherent generative model over the pair $(\boldsymbol{x},\boldsymbol{y})$. In doing so, C-DPS naturally integrates both the learned prior and the forward measurement model into a single, unified framework.
$\bullet$ We propose C-DPS, a novel coupled stochastic framework that introduces a parallel diffusion process in the measurement space $\{\boldsymbol{y}_{t}\}$, evolving jointly with the data-space process $\{\boldsymbol{x}_{t}\}$. This formulation enables principled posterior sampling in diffusion models by treating $(\boldsymbol{x},\boldsymbol{y})$ as a unified generative process.
$\bullet$ By explicitly constructing a Markov chain over $\boldsymbol{y}_{t}$, we derive a closed-form posterior transition $p(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t},\boldsymbol{y}_{t-1})$, eliminating the need to approximate or learn the likelihood term $p(\boldsymbol{y}\mid\boldsymbol{x}_{t})$. This allows for direct integration of the measurement model into the sampling procedure, ensuring consistent Bayesian updates at every diffusion step.
$\bullet$ We develop a scalable and efficient sampling algorithm for C-DPS based on a pre-whitened conjugate gradient solver. This matrix-free implementation retains the runtime efficiency of conventional DPS methods, despite the added complexity of coupled data-measurement diffusion.
$\bullet$ We validate C-DPS through extensive experiments on standard benchmarks, including FFHQ karras2019style and ImageNet deng2009imagenet. Our method achieves state-of-the-art performance across multiple inverse problem settings—such as inpainting, deblurring, and super-resolution—both qualitatively and quantitatively.
Notation: Scalars are represented by non-bold letters, (e.g., $a$ or $A$), vectors by bold lowercase letters (e.g., $\mathbb{a}$), and matrices by bold uppercase letters (e.g., $\mathbb{A}$). The real axis is denoted by $\mathbb{R}$. The symbols $\mathbb{0}$ and $\mathbb{I}$ represent the zero vector and the identity matrix, respectively.
2 Background and Preliminaries
2.1 Diffusion Models
Diffusion models are built upon two key components: a forward noising process and a backward denoising process. These processes operate as Markov chains that progressively perturb and then recover data distributions. In the discrete formulation ddpm, the forward process gradually adds Gaussian noise to the data according to a predefined variance schedule $\{\beta_{t}\}_{t=1}^{T}$, and is described as
$p(\boldsymbol{x}_{1:T}|\boldsymbol{x}_{0})$ $=\prod_{k=1}^{T}p(\boldsymbol{x}_{t}|\boldsymbol{x}_{t-1}),$ (1a)
$p(\boldsymbol{x}_{t}|\boldsymbol{x}_{t-1})$ $=\mathcal{N}(\sqrt{1-\beta_{t}}\boldsymbol{x}_{t-1},\beta_{t}\mathbf{I}),$ (1b)
where $\boldsymbol{x}_{t}\in\mathbb{R}^{d}$ and $\beta_{t}\in(0,1)$ is a monotonically increasing sequence controlling the rate of noise addition over time. Since each transition $p(\boldsymbol{x}_{t}|\boldsymbol{x}_{t-1})$ follows a linear Gaussian model, the marginal distribution $p(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})$ remains Gaussian, given by
$p(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})\sim\mathcal{N}(\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}_{0},(1-\bar{\alpha}_{t})\mathbb{I}),$ (2)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{j=1}^{t}\alpha_{j}$ are derived from the variance schedule ddpm.
To generate data samples, a neural network $\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},t)$ is trained to approximate the score function $\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{x}_{t})$. The reverse denoising process is also formulated as a Markov chain, where the model iteratively reconstructs the data distribution. The backward process is expressed as
$p_{\theta}(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t})$ $=\mathcal{N}\big(\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t}),\boldsymbol{\Sigma}_{\theta}\big),$ (3a)
$\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t})$ $=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\boldsymbol{x}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\boldsymbol{\hat{x}}_{0}(\boldsymbol{x}_{t}),$ (3b)
$\boldsymbol{\Sigma}_{\theta}$ $=\beta_{t}\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\mathbf{I},$ (3c)
where $\boldsymbol{\hat{x}}_{0}(\boldsymbol{x}_{t})$ is the predicted initial state of the data computed using Tweedie’s formula efron2011tweedie:
$\boldsymbol{\hat{x}}_{0}(\boldsymbol{x}_{t})=\frac{\boldsymbol{x}_{t}+(1-\bar{\alpha}_{t})\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}.$ (4)
2.2 Diffusion-Based Inverse Problem Solving
We focus on linear inverse problems, where the goal is to reconstruct an unknown signal $\boldsymbol{x}_{0}\in\mathbb{R}^{d}$ from noisy, incomplete measurements $\boldsymbol{y}\in\mathbb{R}^{m}$, governed by the linear model $\boldsymbol{y}=\mathbb{A}\boldsymbol{x}_{0}+\mathbf{n}$, where $\mathbb{A}\in\mathbb{R}^{m\times d}$ is a known measurement operator, and $\mathbf{n}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}_{\boldsymbol{n}})$ denotes additive Gaussian noise with known covariance $\boldsymbol{\Sigma}_{\boldsymbol{n}}$. This leads to a Gaussian likelihood $p(\boldsymbol{y}|\boldsymbol{x}_{0})=\mathcal{N}(\mathbb{A}\boldsymbol{x}_{0},\boldsymbol{\Sigma}_{\boldsymbol{n}})$.
Recent advances in applying diffusion models to inverse problems can be broadly categorized into two methodological paradigms: projection-based and likelihood-based approaches. These methods are summarized in Algorithm˜1, and a comprehensive discussion is provided in Appendix˜A. We briefly describe each paradigm below to set the stage for our proposed method.
$\bullet$ Projection-based approaches. These approaches iteratively project the generated samples onto the feasible set defined by the measurement $\boldsymbol{y}$. At each reverse diffusion step, the generated sample $\boldsymbol{x}_{t}$ is updated using a projection operator $\mathcal{P}_{\boldsymbol{y}}$, such as
$\boldsymbol{x}_{t}\leftarrow\mathcal{P}_{\boldsymbol{y}}(\boldsymbol{x}_{t}):=\arg\min_{\boldsymbol{z}}\|\boldsymbol{z}-\boldsymbol{x}_{t}\|^{2}\quad\text{subject to}\quad\|\mathbb{A}\boldsymbol{z}-\boldsymbol{y}\|^{2}\leq\epsilon,$
or in simpler forms (e.g., for noiseless measurements),
$\boldsymbol{x}_{t}\leftarrow\boldsymbol{x}_{t}-\mathbb{A}^{\top}(\mathbb{A}\mathbb{A}^{\top})^{-1}(\mathbb{A}\boldsymbol{x}_{t}-\boldsymbol{y}).$
This family includes methods such as ILVR choi2021ilvr, DDRM kawar2022denoising, and DSG ye2024decomposed, which leverage such projections to maintain measurement consistency throughout sampling.
$\bullet$ Likelihood-based approaches. These methods aim to recover the unknown signal $\boldsymbol{x}_{0}$ by approximately sampling from the posterior distribution $p(\boldsymbol{x}_{0}\mid\boldsymbol{y})$, where $\boldsymbol{y}$ denotes the measurement. Using Bayes’ rule, the posterior can be expressed as
$p(\boldsymbol{x}_{0}\mid\boldsymbol{y})\propto p(\boldsymbol{y}\mid\boldsymbol{x}_{0})p(\boldsymbol{x}_{0}).$
Here, $p(\boldsymbol{x}_{0})$ is modeled using a diffusion model trained on clean data, and $p(\boldsymbol{y}\mid\boldsymbol{x}_{0})$ represents the likelihood induced by the forward measurement process (e.g., a linear operator with additive noise).
Since the likelihood is typically intractable to evaluate directly in diffusion models, several methods approximate it or its gradient. For example, DPS DPS and related methods boys2023tweedie; NEURIPS2024_2f46ef57; NEURIPS2024_03261886; hamididiffusion use Tweedie’s formula or Bayesian denoising estimators to approximate $\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{y}\mid\boldsymbol{x}_{t})$ at intermediate timesteps $t$. This approximation is then used to adjust the reverse dynamics by modifying the model’s predicted mean. Specifically, the mean used in the reverse step is replaced by $\tilde{\boldsymbol{\mu}}_{\theta}(\boldsymbol{x}_{t},\boldsymbol{y})=\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t})+\lambda\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{y}\mid\boldsymbol{x}_{t})$, where $\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t})$ is the original predicted mean, and $\lambda$ is a scaling factor.
Algorithm 1 Baseline methods
1:# time steps $T$, $\boldsymbol{y}$, noise schedule $\{\beta_{t}\}$, measurement $\mathbb{A}$, $\{\tilde{\sigma}_{t}\}$.
[…]
3:for $t=T-1,T-2,\dots,0$ do
4:  $\hat{\boldsymbol{s}}\leftarrow\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},t)$
[…]
6:
$\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t})\leftarrow\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\boldsymbol{x}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\boldsymbol{\hat{x}}_{0}(\boldsymbol{x}_{t})$
7:  $\boldsymbol{z}\sim\mathcal{N}(\mathbb{0},\mathbb{I})$.
8:  $\boldsymbol{x}_{t-1}^{\prime}\leftarrow\boldsymbol{\mu}_{\theta}(\boldsymbol{x}_{t})+\tilde{\sigma}_{t}\boldsymbol{z}$
9:  $\boldsymbol{x}_{t-1}$ is obtained from $\boldsymbol{x}_{t-1}^{\prime}$ by either a projection step song2023pseudoinverse; NEURIPS2024_54fa8255; choi2021ilvr; ye2024decomposed; MCG; DSG; kawar2022denoising; hamidi2025conditional or likelihood approximation DPS; boys2023tweedie; NEURIPS2024_2f46ef57; NEURIPS2024_03261886; hamididiffusion.
10:end for
[…]
Algorithm 2 C-DPS (one reverse step, clean form)
1:# steps $T$, measurements $\boldsymbol{y}$, schedule $\{\beta_{t}\}$, operator $\boldsymbol{A}$, noise covariance $\boldsymbol{\Sigma}_{\boldsymbol{n}}$
2:Generate $\{\boldsymbol{y}_{t}\}_{t=0}^{T}$ as in Equation˜5; draw $\boldsymbol{x}_{N}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$
3:for $t=T-1,\,T-2,\,\dots,\,0$ do
4:  $\hat{\boldsymbol{s}}\leftarrow\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},t)$
5:  $\boldsymbol{\Sigma}_{\boldsymbol{y}\mid\boldsymbol{x}}\leftarrow\bar{\alpha}_{t-1}\boldsymbol{\Sigma}_{\boldsymbol{n}}+(1-\bar{\alpha}_{t-1})\boldsymbol{I}$, $c_{t}\leftarrow\tfrac{1-\beta_{t}}{\beta_{t}}$
6:  $\boldsymbol{b}_{t-1}\leftarrow(1-\bar{\alpha}_{t-1})\,\boldsymbol{A}\,\hat{\boldsymbol{s}}$
7:  Define the matrix–free precision operator $\Lambda_{t}\boldsymbol{u}\;\leftarrow\;c_{t}\,\boldsymbol{u}\;+\;\boldsymbol{A}^{\!\top}\boldsymbol{\Sigma}_{\boldsymbol{y}\mid\boldsymbol{x}}^{-1}\boldsymbol{A}\,\boldsymbol{u}$.
8:  Solve for $\boldsymbol{\mu}_{\mathrm{post}}$ with CG
9:  $\boldsymbol{v}\leftarrow\text{PW-CG}(\Lambda_{t},\boldsymbol{A},\boldsymbol{\Sigma}_{\boldsymbol{y}\mid\boldsymbol{x}}^{-1},c_{t})$ (Algorithm˜3)
10:  Update: $\boldsymbol{x}_{t-1}\leftarrow\boldsymbol{\mu}_{\mathrm{post}}+\boldsymbol{v}$
11:end for
[…]
3 Methodology
3.1 Motivation
As discussed in Section˜2.2, most existing approaches to solving inverse problems with diffusion models rely on retro-fitting measurement information into the learned data prior $p(\boldsymbol{x}_{0})$. This is typically done either through heuristic projection-based techniques or by approximating the likelihood $p(\boldsymbol{y}\mid\boldsymbol{x})$. However, both approaches introduce fundamental limitations: projection-based constraints often require manual tuning and lack theoretical justification, while likelihood approximations can distort the posterior geometry, especially under high noise levels (see Figure˜1).
To address these challenges, we propose C-DPS, a principled framework that integrates the measurement process directly into the diffusion dynamics. Central to our idea is an auxiliary forward stochastic process in the measurement space, denoted by $\{\boldsymbol{y}_{t}\}_{t=0}^{T}$. By coupling the data-space process $\{\boldsymbol{x}_{t}\}_{t=0}^{T}$ and $\{\boldsymbol{y}_{t}\}_{t=0}^{T}$, C-DPS enables the derivation of a closed-form posterior distribution $p(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t},\boldsymbol{y}_{t-1})$. This eliminates the need for manually tuning constraint terms or approximating the likelihood. Crucially, unlike the data-space diffusion process which requires training a neural network to approximate reverse dynamics, the measurement-space process $\{\boldsymbol{y}_{t}\}$ is purely forward and analytically defined. It evolves from the observed measurement $\boldsymbol{y}$, not to generate $\boldsymbol{y}_{0}$, but to propagate measurement information coherently across diffusion steps. This coupling allows C-DPS to incorporate observation structure directly into posterior updates, resulting in more stable, interpretable, and accurate reconstructions across a wide range of inverse problems.
It is worth noting that similar measurement-space diffusions have been discussed conceptually in Appendix I.4 of song2020score and further explored in song2022solving for medical imaging, but neither derives an analytic posterior or provides a tractable sampling rule. In contrast, C-DPS formulates a closed-form posterior $p(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t},\boldsymbol{y}_{t-1})$, removing heuristic updates and manual tuning.
In the following subsections, we elaborate how the sequences $\{\boldsymbol{y}_{t}\}_{t=0}^{T}$ and $\{\boldsymbol{x}_{t}\}_{t=0}^{T}$ are generated.
3.2 Constructing a Markov Chain in the Measurement Space
Similarly to the diffusion process in the data space which is defined with Markov chain in Equation˜1b, we define a Markov chain $\{\boldsymbol{y}_{t}\}_{t=0}^{T}$ that starts from the real-world measurement $\boldsymbol{y}_{0}$ and progressively adds Gaussian noise at each forward step. Specifically, we let
$\boldsymbol{y}_{0}$ $\;=\;\mathbb{A}\boldsymbol{x}_{0}+\mathbf{n},$ (5a)
$\boldsymbol{y}_{t}$ $\;=\;\sqrt{1-\beta_{t}}\;\boldsymbol{y}_{t-1}\;+\;\sqrt{\beta_{t}}\;\boldsymbol{z}_{t},$ (5b)
where $\boldsymbol{z}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\beta_{t}\in(0,1)$ determines the noise schedule. The linear operator $\mathbb{A}$ maps the unknown data $\boldsymbol{x}_{0}$ to the measurement space, and $\mathbf{n}$ is the measurement noise.
Remark 1 (Choice of noise schedule).
We use the same noise schedule $\{\beta_{t}\}$ for both the data-space diffusion $\{\boldsymbol{x}_{t}\}$ and the measurement-space diffusion $\{\boldsymbol{y}_{t}\}$. This design choice keeps the two processes synchronized, which simplifies the derivation of the backward update $p(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t},\boldsymbol{y}_{t-1})$ and avoids introducing additional hyperparameters. Conceptually, using a shared schedule ensures that the same fraction of noise is injected at each time step in both domains, preserving a one-to-one correspondence between $\boldsymbol{x}_{t}$ and $\boldsymbol{y}_{t}$. Although it is possible in principle to define separate schedules for the data and measurement spaces, we find that a shared schedule leads to strong empirical performance and a cleaner theoretical formulation.
From the standard forward-diffusion identity, $\boldsymbol{y}_{t}$ could be written in terms of $\boldsymbol{y}_{0}$ as
$\boldsymbol{y}_{t}\;=\;\sqrt{\bar{\alpha}_{t}}\;\,\boldsymbol{y}_{0}\;+\;\sqrt{\,1-\bar{\alpha}_{t}}\;\boldsymbol{\zeta},$ (6)
where $\boldsymbol{\zeta}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\bar{\alpha}_{t}$ is defined in Equation˜2. Further noting $\boldsymbol{y}_{0}=\mathbb{A}\,\mathbf{x}_{0}+\mathbf{n}\;$ with $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\mathbf{n}})$, we have
$\boldsymbol{y}_{t}=\sqrt{\bar{\alpha}_{t}}\,\bigl(\mathbb{A}\,\mathbf{x}_{0}+\mathbf{n}\bigr)\;+\;\sqrt{1-\bar{\alpha}_{t}}\;\boldsymbol{\zeta}.$ (7)
From Equation˜7, it is easy to obtain the following conclusion:
Proposition 1 (Distribution of $\{\boldsymbol{y}_{t}\}$).
For the sequence $\{\boldsymbol{y}_{t}\}_{t=0}^{T}$ defined in Equation˜5, the distribution of $\boldsymbol{y}_{t}$ is a Gaussian whose mean and covariance at step $t$ are given by
$\mathbf{\mu}_{\boldsymbol{y},t}=\;\sqrt{\bar{\alpha}_{t}}\,\mathbb{A}\,\boldsymbol{x}_{0},$ (8a)
$\mathbf{\Sigma}_{\boldsymbol{y},t}=\;\bar{\alpha}_{t}\,\boldsymbol{\Sigma}_{\mathbf{n}}\;+\;\bigl(1-\bar{\alpha}_{t}\bigr)\,\mathbf{I}.$ (8b)
3.3 Generating $\{\boldsymbol{x}_{t}\}$ Consistent with $\{\boldsymbol{y}_{t}\}$
Given the measurement sequence $\{\boldsymbol{y}_{t}\}$, we construct a consistent sequence of latent states $\{\boldsymbol{x}_{t}\}$ that evolves under diffusion while remaining aligned with the observations.
$\bullet$ Initialization. We start by sampling
$\boldsymbol{x}_{T}\;\sim\;\mathcal{N}(\mathbf{0},\mathbf{I}).$
$\bullet$ Backward Recursion. For each $t$ from $T$ down to $1$, we generate $\boldsymbol{x}_{t-1}$ by drawing a sample from $p\bigl(\boldsymbol{x}_{t-1}\big|\boldsymbol{x}_{t},\boldsymbol{y}_{t-1}\bigr)$. Based on the Bayesian update rule for the posterior, we can write [^1] Please refer to Remark 2 to see why the prior $p(\boldsymbol{x}_{t-1})$ is dropped from Equation 9.
$p\bigl(\boldsymbol{x}_{t-1}\,\big|\;\boldsymbol{x}_{t},\boldsymbol{y}_{t-1}\bigr)\;\propto\;p\bigl(\boldsymbol{x}_{t}\,\big|\;\boldsymbol{x}_{t-1}\bigr)\;p\bigl(\boldsymbol{y}_{t-1}\,\big|\;\boldsymbol{x}_{t-1}\bigr).$ (9)
To find $p(\boldsymbol{y}_{t-1}|\boldsymbol{x}_{t-1})$, at an intermediate diffusion step, we use the formula in Equation˜7, and replace $\boldsymbol{x}_{0}$ in this formula with its estimate in terms of $\boldsymbol{x}_{t-1}$ using Equation˜4. Thus, we get
$\boldsymbol{y}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\bigg(\mathbb{A}\bigg[\frac{\boldsymbol{x}_{t-1}+(1-\bar{\alpha}_{t-1})\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t-1},t-1)}{\sqrt{\bar{\alpha}_{t-1}}}\bigg]+\mathbf{n}\bigg)+\sqrt{1-\bar{\alpha}_{t-1}}\;\boldsymbol{\zeta}$ (10)
$=\mathbb{A}\boldsymbol{x}_{t-1}+\mathbb{A}\big((1-\bar{\alpha}_{t-1})\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t-1},t-1)\big)+\sqrt{\bar{\alpha}_{t-1}}\mathbf{n}+\sqrt{1-\bar{\alpha}_{t-1}}\boldsymbol{\zeta}.$ (11)
Since $\mathbf{n}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}_{\boldsymbol{n}})$ and $\boldsymbol{\zeta}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $p(\boldsymbol{y}_{t-1}|\boldsymbol{x}_{t-1})$ is Gaussian with the following parameters:
$p(\boldsymbol{y}_{t-1}|\boldsymbol{x}_{t-1})\sim\mathcal{N}\Bigl(\boldsymbol{\mu}_{\boldsymbol{y}|\boldsymbol{x}},\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}\Bigr),$ (12a)
$\text{where}\qquad\boldsymbol{\mu}_{\boldsymbol{y}|\boldsymbol{x}}=\mathbb{A}\boldsymbol{x}_{t-1}+(1-\bar{\alpha}_{t-1})\mathbb{A}\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t-1},t-1),$ (12b)
$\text{and}\qquad\quad\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}=\bar{\alpha}_{t-1}\boldsymbol{\Sigma}_{\mathbf{n}}+(1-\bar{\alpha}_{t-1})\mathbf{I}.$ (12c)
Substituting Equation˜1b and Equation˜12a into Equation˜9, the posterior $p\bigl(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t},\boldsymbol{y}_{t-1}\bigr)$ could be obtained using Bayes’ rule and dropping the normalizing constant:
$p(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t},\boldsymbol{y}_{t-1})\propto\exp\!\Bigl[-\tfrac{1}{2\beta_{t}}\bigl\|\boldsymbol{x}_{t}-\sqrt{1-\beta_{t}}\,\boldsymbol{x}_{t-1}\bigr\|^{2}-\tfrac{1}{2}\bigl(\boldsymbol{y}_{t-1}-\boldsymbol{\mu}_{\boldsymbol{y}|\boldsymbol{x}}\bigr)^{\!\top}\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}^{-1}\bigl(\boldsymbol{y}_{t-1}-\boldsymbol{\mu}_{\boldsymbol{y}|\boldsymbol{x}}\bigr)\Bigr].$ (13)
The expression in Equation˜13 defines the exact un-normalized posterior density. The first term is quadratic in $\boldsymbol{x}_{t-1}$ and arises from the diffusion prior. However, the second term involves the score network $\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t-1},t-1)$, making the conditional mean $\boldsymbol{\mu}_{\boldsymbol{y}|\boldsymbol{x}}$ a nonlinear function of $\boldsymbol{x}_{t-1}$.
To make the posterior distribution in Equation˜13 Gaussian, we apply a common approximation in diffusion-based inference and freeze the score network at the current iterate:
$\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t-1},t-1)\;\;\longrightarrow\;\;\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},\,t).$ (14)
This substitution avoids evaluating the score at $\boldsymbol{x}_{t-1}$, which is not yet known during sampling, and instead uses the available point $\boldsymbol{x}_{t}$. Intuitively, since $\boldsymbol{x}_{t-1}$ and $\boldsymbol{x}_{t}$ are close for small $\beta_{t}$, this approximation preserves consistency. While this approximation has been used in prior works choi2021ilvr; MCG, we further empirically justify it in the context of our work in Appendix˜B.
Using Equation˜14, $\boldsymbol{\mu}_{\boldsymbol{y}|\boldsymbol{x}}$ becomes affine in $\boldsymbol{x}_{t-1}$,
$\boldsymbol{\mu}_{\boldsymbol{y}|\boldsymbol{x}}=\mathbb{A}\boldsymbol{x}_{t-1}+\underbrace{\!\bigl(1-\bar{\alpha}_{t-1}\bigr)\mathbb{A}\,\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},t)}_{\triangleq\,\boldsymbol{b}_{t-1}},$ (15)
so the product of the two Gaussians in Equation˜9 remains Gaussian $\mathcal{N}\!\bigl(\boldsymbol{\mu}_{\text{post}},\boldsymbol{\Sigma}_{\text{post}}\bigr)$. Collecting quadratic and linear terms gives
$\boldsymbol{\Sigma}_{\text{post}}^{-1}$ $=\frac{1-\beta_{t}}{\beta_{t}}\mathbf{I}+\mathbb{A}^{\!\top}\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}^{-1}\mathbb{A},$ (16a)
$\boldsymbol{\mu}_{\text{post}}$ $=\boldsymbol{\Sigma}_{\text{post}}\Bigl[\frac{\sqrt{1-\beta_{t}}}{\beta_{t}}\,\boldsymbol{x}_{t}+\mathbb{A}^{\!\top}\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}^{-1}\bigl(\boldsymbol{y}_{t-1}-\boldsymbol{b}_{t-1}\bigr)\Bigr].$ (16b)
Hence, C-DPS samples from $\mathcal{N}(\boldsymbol{\mu}_{\text{post}},\boldsymbol{\Sigma}_{\text{post}})$ at each reverse step. The full procedure is shown in Algorithm˜2, with the latent variant, termed LC-DPS, detailed in Appendix˜C. In addition, in Appendix˜D, we extend the above approach for the case of non-linear measurement.
Remark 2.
Note that the exact form of Equation˜9 includes the prior term $p(\boldsymbol{x}_{t-1})$. In diffusion models, after sufficient noising steps under linear or cosine schedules, the marginal $p(\boldsymbol{x}_{t-1})$ is close to $\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ since the noise dominates. Including this term multiplies Equation˜13 by $\exp(-\tfrac{1}{2}\lVert\boldsymbol{x}_{t-1}\rVert_{2}^{2})$, which adds $\boldsymbol{I}$ to the posterior precision, that is $\Lambda_{t}\leftarrow\tfrac{1-\beta_{t}}{\beta_{t}}\boldsymbol{I}+\boldsymbol{A}^{\top}\boldsymbol{\Sigma}_{\boldsymbol{y}\mid\boldsymbol{x}}^{-1}\boldsymbol{A}+\boldsymbol{I}$. Since $\beta_{t}\ll 1$, the extra $\boldsymbol{I}$ changes each diagonal by at most a small fraction $O(\beta_{t})$, which is numerically negligible. We therefore omit $p(\boldsymbol{x}_{t-1})$ in Equation˜9 for clarity.
3.4 Efficient sampling
Our goal is to draw $\boldsymbol{x}_{t-1}\sim p(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t},\boldsymbol{y}_{t-1})$ as in Equation˜16 without forming dense factorizations. Define the posterior precision operator
$\Lambda_{t}\;=\;\boldsymbol{\Sigma}_{\text{post}}^{-1}\;=\;c_{t}\,\boldsymbol{I}\;+\;\boldsymbol{A}^{\!\top}\boldsymbol{\Sigma}_{\boldsymbol{y}\mid\boldsymbol{x}}^{-1}\boldsymbol{A},\qquad c_{t}\;=\;\tfrac{1-\beta_{t}}{\beta_{t}}.$ (17)
Direct Cholesky on dense $\boldsymbol{\Sigma}_{\boldsymbol{y}\mid\boldsymbol{x}}$ is $\mathcal{O}(d^{3})$ and impractical. We therefore use two matrix-free conjugate-gradient (CG) solves per reverse step, implemented by Algorithms˜2 and 3.
Step 1: mean solve.
Compute $\boldsymbol{\mu}_{\text{post}}$ by solving
$\Lambda_{t}\,\boldsymbol{\mu}_{\text{post}}\;=\;c_{t}\,\boldsymbol{x}_{t}\;+\;\boldsymbol{A}^{\!\top}\boldsymbol{\Sigma}_{\boldsymbol{y}\mid\boldsymbol{x}}^{-1}\bigl(\boldsymbol{y}_{t-1}-\boldsymbol{b}_{t-1}\bigr),\qquad\boldsymbol{b}_{t-1}\,=\,(1-\bar{\alpha}_{t-1})\,\boldsymbol{A}\,\hat{\boldsymbol{s}}$, (18)
with $\hat{\boldsymbol{s}}=\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},t)$.
Step 2: noise draw (PW-CG).
Draw $\boldsymbol{v}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}_{\text{post}})$ by solving $\Lambda_{t}\boldsymbol{v}=\boldsymbol{z}$ for a synthetic right-hand side $\boldsymbol{z}$ satisfying $\operatorname{cov}(\boldsymbol{z})=\Lambda_{t}$. Algorithm 3 shows how to build such a $\boldsymbol{z}$ from two standard Gaussians using a whitening operator.
Step 3: update.
Set
$\boldsymbol{x}_{t-1}\;=\;\boldsymbol{\mu}_{\text{post}}\;+\;\boldsymbol{v}.$ (19)
Pre-whitened conjugate gradient (PW-CG).
Let $\boldsymbol{W}$ satisfy $\boldsymbol{W}^{\!\top}\boldsymbol{W}=\boldsymbol{\Sigma}_{\boldsymbol{y}\mid\boldsymbol{x}}^{-1}$. For many structured noise models a closed form exists. For example, if $\boldsymbol{\Sigma}_{\boldsymbol{n}}=\sigma^{2}\boldsymbol{I}$, then
$\boldsymbol{\Sigma}_{\boldsymbol{y}\mid\boldsymbol{x}}\,=\,\bigl(\bar{\alpha}_{t-1}\sigma^{2}+1-\bar{\alpha}_{t-1}\bigr)\boldsymbol{I}\,\triangleq\,\gamma_{t}\boldsymbol{I},\qquad\boldsymbol{W}\,=\,\gamma_{t}^{-1/2}\boldsymbol{I}.$ (20)
Define $\tilde{\boldsymbol{A}}=\boldsymbol{W}\boldsymbol{A}$ so that $\Lambda_{t}=c_{t}\boldsymbol{I}+\tilde{\boldsymbol{A}}^{\!\top}\tilde{\boldsymbol{A}}$. PW-CG samples
$\boldsymbol{\varepsilon}_{1}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{d}),\qquad\boldsymbol{\varepsilon}_{2}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{m}),\qquad\boldsymbol{z}\,=\,\sqrt{c_{t}}\,\boldsymbol{\varepsilon}_{1}\,+\,\tilde{\boldsymbol{A}}^{\!\top}\boldsymbol{\varepsilon}_{2},$ (21)
which gives $\operatorname{cov}(\boldsymbol{z})=\Lambda_{t}$. Solving $\Lambda_{t}\boldsymbol{v}=\boldsymbol{z}$ by CG then yields $\boldsymbol{v}\sim\mathcal{N}(\boldsymbol{0},\Lambda_{t}^{-1})=\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}_{\text{post}})$.
Practicalities and cost.
Each CG iteration applies $\boldsymbol{A}$ and $\boldsymbol{A}^{\!\top}$ once. A diagonal preconditioner $\boldsymbol{P}_{t}=\operatorname{diag}(\Lambda_{t})$ works well. CG requires $\mathcal{O}(\sqrt{\kappa})$ iterations with $\kappa$ the condition number of $\boldsymbol{P}_{t}^{-1}\Lambda_{t}$. Empirically $\kappa<50$ across our tasks, so the PW-CG cost per step is $\mathcal{O}\bigl(\sqrt{\kappa}\cdot\mathrm{nnz}(\boldsymbol{A})\bigr)$. In our setup, score evaluation dominates the runtime, while the linear solves add a small overhead.
To contextualize this cost, we note that evaluating the score network $\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},t)$—typically implemented as a U-Net or Transformer—takes approximately 100–200 milliseconds per step on a modern GPU (e.g., NVIDIA V100 or P100) for a $256\times 256$ input. In contrast, our linear solver runs significantly faster, often requiring under 10 milliseconds. Its contribution to the total runtime is therefore negligible compared with the dominant cost of score evaluation.
Table 1: Quantitative results on the $1\mathrm{k}$ validation sets of FFHQ $256\times 256$ and ImageNet $256\times 256$. Bold and underline indicate the best and second-best results, respectively. Green and red denote performance improvements and degradations relative to the best baseline.
Pixel-Domain Methods
Dataset Method Inpaint (Random) Inpaint (Box) Deblur (Gaussian) Deblur (Motion) SR ($4\times$)
FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
FFHQ DPS 21.19 0.212 0.851 33.12 0.168 0.873 44.05 0.257 0.811 39.92 0.242 0.859 39.35 0.214 0.852
[…]
DDRM 69.71 0.587 0.319 42.93 0.204 0.869 74.92 0.332 0.767 - - - 62.15 0.294 0.835
MCG 29.26 0.286 0.751 40.11 0.309 0.703 101.2 0.340 0.051 - - - 87.64 0.520 0.559
ILVR 25.74 0.231 0.672 37.24 0.175 0.854 52.93 0.297 0.784 - - - 47.59 0.253 0.844
ReSample 21.25 0.202 0.847 33.51 0.160 0.866 37.05 0.251 0.822 31.19 0.220 0.892 30.48 0.204 0.851
PnP-ADMM 123.6 0.692 0.325 151.9 0.406 0.642 90.42 0.441 0.812 - - - 66.52 0.353 0.855
Score-SDE 76.54 0.612 0.437 60.06 0.331 0.678 109.0 0.403 0.109 - - - 96.72 0.563 0.617
ADMM-TV 181.5 0.463 0.784 68.94 0.322 0.814 186.7 0.507 0.801 - - - 110.6 0.428 0.803
PnP-DM 21.15 0.208 0.858 32.21 0.155 0.877 41.92 0.251 0.816 37.21 0.233 0.871 36.21 0.210 0.859
DAPS 20.77 0.201 0.869 29.44 0.144 0.882 35.84 0.242 0.830 30.26 0.215 0.911 30.15 0.202 0.854
DMPlug 20.12 0.197 0.877 27.12 0.140 0.888 32.44 0.230 0.830 27.55 0.210 0.925 28.55 0.199 0.862
C-DPS 20.14 0.195 0.881 26.33 0.132 0.871 32.24 0.238 0.832 27.29 0.217 0.921 28.41 0.196 0.855
ImageNet DPS 35.87 0.303 0.739 38.82 0.262 0.794 62.72 0.444 0.706 56.08 0.389 0.634 50.66 0.337 0.781
[…]
DDRM 114.9 0.665 0.403 45.95 0.245 0.814 63.02 0.427 0.705 - - - 59.57 0.339 0.790
MCG 39.19 0.414 0.546 39.74 0.330 0.633 95.04 0.550 0.441 - - - 144.5 0.637 0.227
ILVR 38.27 0.372 0.656 39.51 0.278 0.726 71.24 0.421 0.662 - - - 95.3 0.532 0.498
ReSample 33.47 0.289 0.730 39.54 0.259 0.799 61.24 0.439 0.708 55.76 0.370 0.637 49.19 0.339 0.777
PnP-ADMM 114.7 0.677 0.300 78.24 0.367 0.657 100.6 0.519 0.669 - - - 97.27 0.433 0.761
Score-SDE 127.1 0.659 0.517 54.07 0.354 0.612 120.3 0.667 0.436 - - - 170.7 0.701 0.256
ADMM-TV 189.3 0.510 0.676 87.69 0.319 0.785 155.7 0.588 0.634 - - - 130.9 0.523 0.679
PnP-DM 34.92 0.296 0.736 37.67 0.258 0.797 61.06 0.433 0.707 55.33 0.372 0.636 50.10 0.336 0.786
DAPS 33.94 0.282 0.741 35.46 0.248 0.801 60.12 0.419 0.709 54.82 0.365 0.639 49.62 0.333 0.789
DMPlug 32.85 0.226 0.748 34.28 0.247 0.804 57.42 0.407 0.714 53.13 0.366 0.642 48.96 0.324 0.793
C-DPS 32.37 0.214 0.755 33.24 0.236 0.807 56.36 0.391 0.712 52.06 0.352 0.644 47.30 0.316 0.795
Latent-Domain Methods
Dataset Method Inpaint (Random) Inpaint (Box) Deblur (Gaussian) Deblur (Motion) SR ($4\times$)
FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$
FFHQ PSLD 47.21 0.221 0.809 43.02 0.158 0.813 89.51 0.316 0.631 96.15 0.336 0.678 74.36 0.287 0.649
ReSample 39.85 0.140 0.746 53.21 0.184 0.749 71.69 0.255 0.714 44.72 0.198 0.823 93.18 0.392 0.594
RLSD 38.25 0.142 0.808 44.08 0.153 0.812 68.92 0.244 0.750 49.10 0.284 0.810 61.37 0.203 0.774
LC-DPS 36.67 0.137 0.815 42.11 0.144 0.821 65.71 0.232 0.759 46.57 0.272 0.819 58.41 0.197 0.787
ImageNet PSLD 83.21 0.337 0.783 146.53 0.465 0.694 91.39 0.390 0.688 124.67 0.511 0.594 97.45 0.360 0.694
ReSample 59.87 0.143 0.756 127.84 0.262 0.631 65.35 0.254 0.703 66.89 0.227 0.738 113.42 0.370 0.576
RLSD 60.44 0.141 0.787 130.16 0.259 0.705 67.34 0.249 0.712 68.43 0.236 0.735 92.73 0.317 0.680
LC-DPS 60.02 0.140 0.790 113.67 0.252 0.714 67.21 0.244 0.718 63.94 0.216 0.748 88.15 0.288 0.683
Algorithm 3 PW-CG (draw $\boldsymbol{v}\!\sim\!\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}_{\mathrm{post}})$ without Cholesky)
1:precision operator $\Lambda_{t}$, matrix $\boldsymbol{A}$, action of $\boldsymbol{\Sigma}_{\boldsymbol{y}\mid\boldsymbol{x}}^{-1}$ (or its square root), scalar $c_{t}$
2:Draw $\boldsymbol{\varepsilon}_{1}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{d})$, $\boldsymbol{\varepsilon}_{2}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{m})$, independent
3:(Prewhiten) define a whitening operator $\boldsymbol{W}$ with $\boldsymbol{W}^{\!\top}\boldsymbol{W}=\boldsymbol{\Sigma}_{\boldsymbol{y}\mid\boldsymbol{x}}^{-1}$, and set $\tilde{\boldsymbol{A}}=\boldsymbol{W}\boldsymbol{A}$
4:Form $\boldsymbol{z}\leftarrow\sqrt{c_{t}}\,\boldsymbol{\varepsilon}_{1}+\tilde{\boldsymbol{A}}^{\!\top}\boldsymbol{\varepsilon}_{2}$
5:Solve $\boldsymbol{v}\leftarrow\text{CG-solve}(\Lambda_{t},\boldsymbol{z})$ (that is, solve $\Lambda_{t}\boldsymbol{v}=\boldsymbol{z}$)
[…]
4 Experiments
4.1 Quantitative Results
$\bullet$ Experimental setup. Following prior work DPS; dou2024diffusion, we evaluate our method on the FFHQ $256{\times}256$ karras2019style and ImageNet $256{\times}256$ deng2009imagenet datasets, using 1,000 validation images from each. All images are normalized to the $[0,1]$ range. To ensure a fair comparison, we adopt the same experimental settings used in DPS across all evaluated methods. The measurement data is corrupted with additive Gaussian noise of zero mean and standard deviation $\sigma=0.05$. During inference, we use a fixed number of reverse diffusion steps $T=1000$, following standard practice in the literature. For score estimation, we utilize the pre-trained model from DPS for FFHQ and the model from dhariwal2021diffusion for ImageNet.
The measurement models used in our experiments follow those described in DPS: ($i$) Inpainting: For box-type inpainting, a central $128\times 128$ region is masked out; for random-type inpainting, 92% of pixels (across all RGB channels) are randomly masked. ($ii$) Super-resolution (SR): Low-resolution measurements are obtained by applying bicubic downsampling. (iiiiii) Gaussian blur: A Gaussian kernel of size $61\times 61$ with a standard deviation of 5.0 is convolved with the image. ($iv$) Motion blur: Motion blur kernels of size $61\times 61$ are generated using the code from motionblur, with an intensity parameter of 0.5. These kernels are convolved with the ground-truth images to produce the measurements.
[…]
$\bullet$ Evaluation metrics. To evaluate the performance of different methods, we adopt the metrics used in DPS: ($i$) learned perceptual image patch similarity (LPIPS) zhang2018unreasonable, ($ii$) Frechet inception distance (FID) heusel2017gans, and (iiiiii) structure similarity index measure (SSIM) wang2004image. In addition, we provide results for peak signal-to-noise ratio (PSNR) in Appendix˜F. All experiments are conducted using a single NVIDIA P100 GPU with 12 GB of memory.
$\bullet$ Results. The quantitative results for both datasets are presented in Table˜1. Across nearly all tasks, C-DPS (or LC-DPS) outperforms the baseline methods. In a few settings, DMPlug attains scores that are competitive with ours. We emphasize that although DMPlug can be close in accuracy, its runtime is approximately $4\times$ slower than C-DPS, which limits its practical utility; see Section˜F.1 for run-time details. We also conduct a similar set of experiments under varying Gaussian noise levels, presented in Section˜H.1.
Figure 2: Qualitative results on FFHQ dataset.
4.2 Visual Comparison
In this subsection, we provide a visual comparison of reconstructions produced by C-DPS, DPS, and ReSample. We choose to include DPS and ReSample in these visual comparisons because, as shown in Table 1, they achieve the strongest quantitative performance among the evaluated baselines. We randomly select five images from the FFHQ test set and corrupt them using the measurement models described in Section˜4.1, with minor adjustments to enhance visual clarity: for super-resolution, we apply an $8\times$ downsampling factor, and for random inpainting, 95% of the pixels are masked. The reconstructed images are presented in Figure˜2, where each row corresponds to a different measurement setting. More qualitative results are presented in Section˜H.2.
As observed, C-DPS consistently produces reconstructions that more closely resemble the ground truth, exhibiting fewer artifacts and better perceptual quality compared with DPS and ReSample.
4.3 Measurement Fidelity
In this subsection, we demonstrate that C-DPS achieves higher measurement fidelity compared with DPS and ReSample. Specifically, we show that C-DPS recovers solutions $\boldsymbol{x}$ that better satisfy the measurement model.
To this end, we conduct a motion deblurring experiment under a noiseless setting, where the Gaussian noise is removed (i.e., $\boldsymbol{y}=\mathbb{A}\boldsymbol{x}_{0}$). We then track the reconstruction error $\|\boldsymbol{y}-\mathbb{A}\boldsymbol{x}_{t}\|_{2}^{2}$ as a function of sampling progress. The progress is expressed as a percentage relative to the total number of reverse diffusion steps taken by each method.
We run each method over 100 instances and report the average error along with the shaded region representing the range of observed values. The results are shown in Figure˜4. As illustrated, C-DPS consistently maintains lower measurement error throughout the sampling trajectory, indicating better alignment with the forward measurement model.
4.4 Why C-DPS Works: Posterior Recovery on a Ground-Truth Benchmark
In this subsection, we aim to demonstrate that the superior performance of C-DPS stems from its improved approximation of the true posterior distribution. To support this claim, we compare its ability to recover the posterior against the best benchmarks DPS and ReSample.
To this end, we construct a toy dataset where the data distribution $p(\boldsymbol{x}_{0})$ is defined as a mixture of 25 Gaussian components. [^2] We follow the dataset construction procedure in cardoso2023monte; boys2023tweedie. The means and variances of the mixture components are detailed in Appendix˜G, where we also explain how the ground-truth posterior can be computed in closed form for any given observation $\boldsymbol{y}$, measurement matrix $\mathbb{A}$, and noise level $\sigma$.
To systematically evaluate each method, we generate measurement models $(\boldsymbol{y},\mathbb{A})\in\mathbb{R}^{m}\times\mathbb{R}^{m\times d}$ across combinations of signal dimension, number of measurements, and noise levels. Specifically, we consider $(d,m,\sigma)\in\{8,80,800\}\times\{1,2,4\}\times\{10^{-2},10^{-1},10^{0}\}$, yielding 27 distinct settings. All Gaussian mixture components are equally weighted to ensure a balanced posterior.
Figure 3: Visualizing the first two dimensions of the estimated posterior distributions for the configuration ($d=80$, $m=1$, $\sigma=10^{-1}$) for a random $\mathbb{A}$.
For each configuration, we generate 1000 samples from the true posterior and apply C-DPS, DPS and ReSample to approximate the posterior using 1000 steps. We then assess the quality of each approximation using the sliced Wasserstein (SW) distance kolouri2019generalized, a metric that captures high-dimensional differences. The SW distance is computed using $10^{4}$ random projections per method.
Table˜2 reports the mean SW distances along with 95% confidence intervals, calculated over 20 randomly sampled measurement matrices $\mathbb{A}$ per configuration. Additionally, Figure˜3 visualizes the estimated posterior in the first two dimensions for the setting $(d=80,m=1,\sigma=0.1)$, using a single randomly drawn measurement matrix. As illustrated, C-DPS more accurately recovers the true posterior, successfully capturing all mixture modes. In contrast, DPS and ReSample either miss certain modes or fail to represent the full posterior geometry, highlighting the advantage of our coupled diffusion formulation.
Figure 4: Evolution of the measurement error $\|\boldsymbol{y}-\mathbb{A}\boldsymbol{x}_{t}\|_{2}^{2}$ over the sampling trajectory for DPS, ReSample, and C-DPS on the FFHQ (left) and ImageNet (right) datasets.
Table 2: Sliced Wasserstein (SW) distance between the true and estimated posterior distributions across different dimensions $d$, numbers of measurements $m$, and noise levels $\sigma$. Lower is better.
$d=8$ $d=80$ $d=800$
Method $m=1$ $m=2$ $m=4$ $m=1$ $m=2$ $m=4$ $m=1$ $m=2$ $m=4$
$\sigma=10^{-2}$ C-DPS 2.2 1.5 0.5 2.9 1.7 0.4 3.3 2.5 0.3
DPS 4.7 1.8 0.7 5.6 3.2 1.2 5.8 3.5 1.4
ReSample 2.6 2.1 3.8 3.2 2.8 0.6 3.5 3.1 0.4
$\sigma=10^{-1}$ C-DPS 1.8 0.9 0.6 2.5 1.7 0.4 2.8 2.3 0.4
DPS 4.7 1.5 0.8 5.1 3.1 1.0 5.7 3.1 1.3
ReSample 2.2 1.6 3.8 2.9 2.7 0.6 3.3 2.7 0.4
$\sigma=10^{0}$ C-DPS 1.2 1.9 0.9 1.7 1.2 0.8 1.6 1.5 0.7
DPS 5.2 3.5 2.5 6.9 3.9 1.7 6.8 4.7 0.9
ReSample 1.5 2.3 1.8 1.6 1.4 0.9 2.0 2.0 0.6
5 Conclusion
We have introduced coupled data and measurement space diffusion posterior sampling (C-DPS), a novel framework that addresses the limitations of likelihood approximation in diffusion-based inverse problem solvers. By introducing a forward stochastic process in the measurement space $\{\boldsymbol{y}_{t}\}$, which evolves in parallel with the data-space diffusion $\{\boldsymbol{x}_{t}\}$, C-DPS enables the derivation of a closed-form posterior distribution. This coupling allows for accurate, recursive sampling without relying on approximations of the likelihood or heuristic constraints. Empirical results demonstrate that C-DPS consistently outperforms state-of-the-art baselines, delivering robust and high-fidelity reconstructions across a wide range of inverse problems. We also discuss limitations and failure cases in Appendix˜I, highlighting opportunities for future improvement.
Acknowledgments
This work was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC).
Appendix A Related Works
In many practical applications, we often encounter the underdetermined regime where $m<d$, making the inverse problem ill-posed. To obtain meaningful reconstructions in such settings, incorporating prior knowledge is essential. Within the Bayesian framework, this is addressed by modeling a prior $p(\boldsymbol{x}_{0})$ and forming the posterior $p(\boldsymbol{x}_{0}|\boldsymbol{y})$ via Bayes’ rule:
$p(\boldsymbol{x}_{0}|\boldsymbol{y})\propto p(\boldsymbol{y}|\boldsymbol{x}_{0})p(\boldsymbol{x}_{0}).$
When employing diffusion models as priors, one can extend the reverse-time stochastic differential euqtion (SDE) used in unconditional models,
$d\boldsymbol{x}=\left[-\tfrac{\beta(t)}{2}\boldsymbol{x}-\beta(t)\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t})\right]dt+\sqrt{\beta(t)}d\bar{\boldsymbol{w}},$
by incorporating an additional likelihood gradient to perform posterior sampling:
$d\boldsymbol{x}=\left[-\tfrac{\beta(t)}{2}\boldsymbol{x}-\beta(t)\left(\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t})+\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{y}|\boldsymbol{x}_{t})\right)\right]dt+\sqrt{\beta(t)}d\bar{\boldsymbol{w}}.$ (22)
This follows directly from the identity:
$\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t}|\boldsymbol{y})=\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t})+\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{y}|\boldsymbol{x}_{t}).$
In this formulation, two quantities are needed: the prior score $\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t})$, which can be obtained from a pre-trained diffusion model, and the likelihood score $\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{y}|\boldsymbol{x}_{t})$. The latter is generally intractable, since the data $\boldsymbol{y}$ is conditionally dependent only on $\boldsymbol{x}_{0}$, not directly on $\boldsymbol{x}_{t}$. As a result, estimating this likelihood gradient becomes a key challenge. In the following sections, we explore various strategies for approximating this term.
Score-ALD [mri_paper] and Score-SDE [song2020score].
Among the earliest approaches for solving linear inverse problems with diffusion models are Score-ALD and Score-SDE, both of which estimate the gradient of the log-likelihood to steer the reverse diffusion trajectory. The key difference lies in how they compute the residual. Score-ALD uses a deterministic correction:
$\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{y}|\boldsymbol{x}_{t})\approx-\frac{\mathbb{A}^{\top}(\boldsymbol{y}-\mathbb{A}\boldsymbol{x}_{t})}{\sigma_{\boldsymbol{y}}^{2}+\gamma_{t}^{2}},$
where $\gamma_{t}$ is a tunable parameter regulating the guidance strength. On the other hand, Score-SDE adds Gaussian noise to the measurements before evaluating the discrepancy:
$\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{y}|\boldsymbol{x}_{t})\approx-\mathbb{A}^{\top}(\boldsymbol{y}+\sigma_{t}\boldsymbol{\epsilon}-\mathbb{A}\boldsymbol{x}_{t}),\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},I_{m}).$
While both methods use the measurement error to guide the reverse process, Score-SDE introduces stochasticity by perturbing the observations, effectively pushing the samples toward a noisy target $\boldsymbol{y}_{t}=\boldsymbol{y}+\sigma_{t}\boldsymbol{\epsilon}$.
ILVR [choi2021ilvr].
ILVR, originally introduced for super-resolution tasks, applies a similar principle by guiding the reverse process with a noised version of the measurements. Its gradient approximation takes the form
$\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{y}|\boldsymbol{x}_{t})\approx-\mathbb{A}^{\dagger}(\boldsymbol{y}_{t}-\mathbb{A}\boldsymbol{x}_{t})=-(\mathbb{A}^{\top}\mathbb{A})^{-1}\mathbb{A}^{\top}(\boldsymbol{y}_{t}-\mathbb{A}\boldsymbol{x}_{t}),$
where $\mathbb{A}^{\dagger}$ denotes the Moore-Penrose pseudo-inverse of $A$, and $\boldsymbol{y}_{t}=\boldsymbol{y}+\sigma_{t}\boldsymbol{\epsilon}$. Compared with Score-SDE, ILVR can be viewed as a preconditioned variant, replacing the simple adjoint $\mathbb{A}^{\top}$ with a full pseudo-inverse to achieve a more accurate projection onto the measurement-consistent space.
DPS [DPS].
While the previous methods are tailored for linear inverse problems, DPS extends to non-linear settings and is among the most widely used reconstruction techniques in this domain. The core approximation behind DPS is
$\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{y}|\boldsymbol{x}_{t})\approx\nabla_{\boldsymbol{x}_{t}}\log p\left(\boldsymbol{y}\mid\boldsymbol{x}_{0}=\mathbb{E}[\boldsymbol{x}_{0}\mid\boldsymbol{x}_{t}]\right).$
Assuming additive Gaussian noise and a forward operator $\mathbb{A}$, the likelihood becomes
$p\left(\boldsymbol{y}\mid\boldsymbol{x}_{0}=\mathbb{E}[\boldsymbol{x}_{0}\mid\boldsymbol{x}_{t}]\right)=\mathcal{N}\left(\boldsymbol{y};\mathbb{A}(\mathbb{E}[\boldsymbol{x}_{0}\mid\boldsymbol{x}_{t}]),\sigma_{\boldsymbol{y}}^{2}I\right),$
yielding the approximation
$\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{y}|\boldsymbol{x}_{t})\approx\frac{1}{\sigma_{\boldsymbol{y}}^{2}}\nabla_{\boldsymbol{x}_{t}}^{\top}\mathbb{A}(\mathbb{E}[\boldsymbol{x}_{0}\mid\boldsymbol{x}_{t}])\left(\mathbb{A}(\mathbb{E}[\boldsymbol{x}_{0}\mid\boldsymbol{x}_{t}])-\boldsymbol{y}\right).$
For linear inverse problems with $\mathbb{A}(\boldsymbol{x})=\mathbb{A}\boldsymbol{x}$, this simplifies to
$\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{y}|\boldsymbol{x}_{t})\approx-\frac{1}{\sigma_{\boldsymbol{y}}^{2}}\nabla_{\boldsymbol{x}_{t}}^{\top}\mathbb{E}[\boldsymbol{x}_{0}\mid\boldsymbol{x}_{t}]\mathbb{A}^{\top}\left(\boldsymbol{y}-A\mathbb{E}[\boldsymbol{x}_{0}\mid\boldsymbol{x}_{t}]\right).$
Using Tweedie’s formula, the gradient term can be further written as
$\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{y}|\boldsymbol{x}_{t})\approx-\frac{1}{\sigma_{\boldsymbol{y}}^{2}}\left(I+\nabla_{\boldsymbol{x}_{t}}^{2}\log p_{t}(\boldsymbol{x}_{t})\right)^{\top}\mathbb{A}^{\top}\left(\boldsymbol{y}-A\mathbb{E}[\boldsymbol{x}_{0}\mid\boldsymbol{x}_{t}]\right).$
In practice, DPS does not rely on the theoretical guidance strength above, and instead employs an adaptive step size inversely scaled by the norm of the measurement error.
MCG [MCG], DSG [DSG], and MPGD [he2024manifold].
Several recent works build on DPS by incorporating geometric insights to improve reconstruction quality. MCG [MCG] provides a geometric interpretation of DPS, demonstrating that its approximation helps maintain samples on the data manifold. Building on this, DSG [DSG] introduces a theoretically grounded step size derived from the MCG perspective, and combines it with projected gradient descent to enhance sample fidelity. More recently, MPGD [he2024manifold] further improves performance by constraining the update steps to lie within a learned low-dimensional subspace using autoencoding, effectively regularizing the sampling path to remain close to the underlying manifold.
[…]
The DPS approximation effectively replaces the posterior $p(\boldsymbol{x}_{0}|\boldsymbol{x}_{t})$ with a Dirac delta centered at its mean:
$p(\boldsymbol{x}_{0}|\boldsymbol{x}_{t})\approx\delta\left(\boldsymbol{x}_{0}-\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]\right).$
[…]
$p(\boldsymbol{x}_{0}|\boldsymbol{x}_{t})\approx\mathcal{N}\left(\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}],r_{t}^{2}I_{n}\right),$
where $r_{t}$ is a tunable hyperparameter. For linear inverse problems, this leads to the marginal likelihood
$p(\boldsymbol{y}|\boldsymbol{x}_{t})\approx\mathcal{N}\left(\mathbb{A}\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}],r_{t}^{2}\mathbb{A}\mathbb{A}^{\top}+\sigma_{\boldsymbol{y}}^{2}I\right),$
and the corresponding gradient becomes
$\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{y}|\boldsymbol{x}_{t})\approx-\frac{\partial\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]}{\partial\boldsymbol{x}_{t}}(r_{t}^{2}\mathbb{A}\mathbb{A}^{\top}+\sigma_{\boldsymbol{y}}^{2}I)^{-1}\mathbb{A}^{\top}\left(\boldsymbol{y}-\mathbb{A}\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]\right).$
This formulation softens the delta approximation of DPS and introduces a covariance-aware correction, improving flexibility in modeling uncertainty.
Moment Matching [rozet2024learning].
[…]
$V[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]=\sigma_{t}^{4}H(\log p_{t}(\boldsymbol{x}_{t}))+\sigma_{t}^{2}I_{n}=\sigma_{t}^{2}\nabla_{\boldsymbol{x}_{t}}\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}],$
resulting in an anisotropic Gaussian approximation:
$p(\boldsymbol{x}_{0}|\boldsymbol{x}_{t})\approx\mathcal{N}(\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}],V[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]).$
For linear inverse problems, this yields a refined estimate for the measurement score:
$\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{y}|\boldsymbol{x}_{t})\approx-\nabla_{\boldsymbol{x}_{t}}\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]^{\top}\mathbb{A}^{\top}\left(\sigma_{\boldsymbol{y}}^{2}I+\mathbb{A}\sigma_{t}^{2}\nabla_{\boldsymbol{x}_{t}}\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]\mathbb{A}^{\top}\right)^{-1}(\boldsymbol{y}-\mathbb{A}\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]).$
In high-dimensional settings, explicitly computing the full Jacobian $\nabla_{\boldsymbol{x}_{t}}\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]$ is computationally prohibitive. To address this, the authors employ automatic differentiation to evaluate Jacobian-vector products efficiently, avoiding the need to form the full matrix.
SNIPS [kawar2021snips] and DDRM [kawar2022denoising].
These methods reformulate linear inverse problems as noisy inpainting in the spectral domain via the singular value decomposition (SVD) of the measurement matrix, $\mathbb{A}=U\Sigma V^{\top}$. With this, the measurement model becomes
$\bar{\boldsymbol{y}}=\Sigma\bar{\boldsymbol{x}}+\sigma_{\boldsymbol{y}}\bar{\boldsymbol{z}},\quad\text{where}\quad\bar{\boldsymbol{x}}=V^{\top}\boldsymbol{x},\,\bar{\boldsymbol{y}}=U^{\top}\boldsymbol{y},\,\bar{\boldsymbol{z}}=U^{\top}\boldsymbol{z}.$
SNIPS solves the inverse problem in this transformed space by performing annealed Langevin dynamics to sample from $p(\bar{\boldsymbol{x}}|\bar{\boldsymbol{y}})$, using the approximation
$\nabla_{\bar{\boldsymbol{x}}_{t}}\log p(\bar{\boldsymbol{y}}|\bar{\boldsymbol{x}}_{t})\approx-\Sigma^{\top}\left|\sigma_{\boldsymbol{y}}^{2}I-\sigma_{t}^{2}\Sigma\Sigma^{\top}\right|^{\dagger}(\bar{\boldsymbol{y}}-\Sigma\bar{\boldsymbol{x}}_{t}).$
DDRM improves upon SNIPS by replacing the noisy iterate $\bar{\boldsymbol{x}}_{t}$ with the posterior mean $\bar{\boldsymbol{x}}_{0|t}=V^{\top}\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]$ in the above expression. Furthermore, DDRM introduces a sampling rule based on conditional Gaussian distributions over the spectral components $\bar{\boldsymbol{x}}_{t}^{(i)}$, which adapt based on the singular value $s_{i}$, the diffusion noise level $\sigma_{t}$, and a hyperparameter $\eta\in(0,1]$ to control stochasticity:
$\bar{\boldsymbol{x}}_{t}^{(i)}\sim\begin{cases}\mathcal{N}(\bar{\boldsymbol{x}}_{0|t+1}^{(i)}+\sqrt{1-\eta^{2}}\sigma_{t}\frac{\bar{\boldsymbol{x}}_{t+1}^{(i)}-\bar{\boldsymbol{x}}_{0|t+1}^{(i)}}{\sigma_{t+1}},\eta^{2}\sigma_{t}^{2}),&s_{i}=0\\ \mathcal{N}(\bar{\boldsymbol{x}}_{0|t+1}^{(i)}+\sqrt{1-\eta^{2}}\sigma_{t}\frac{\bar{\boldsymbol{y}}^{(i)}-\bar{\boldsymbol{x}}_{0|t+1}^{(i)}}{\sigma_{\boldsymbol{y}}/s_{i}},\eta^{2}\sigma_{t}^{2}),&\sigma_{t}<\frac{\sigma_{\boldsymbol{y}}}{s_{i}}\\ \mathcal{N}(\bar{\boldsymbol{y}}^{(i)},\sigma_{t}^{2}-\frac{\sigma_{\boldsymbol{y}}^{2}}{s_{i}^{2}}),&\sigma_{t}\geq\frac{\sigma_{\boldsymbol{y}}}{s_{i}}\end{cases}.$
When $\eta=1$, this sampling reduces to the deterministic posterior formulation used in the original DDRM algorithm.
DDS [chungdecomposed] and DiffPIR [zhu2023denoising].
Both DDS and DiffPIR approximate the posterior mean via a proximal formulation:
$\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t},\boldsymbol{y}]\approx\arg\min_{\boldsymbol{x}}\frac{1}{2}\|\boldsymbol{y}-\mathbb{A}\boldsymbol{x}\|^{2}+\frac{\lambda_{t}}{2}\|\boldsymbol{x}-\mathbb{E}[\boldsymbol{x}_{0}|\boldsymbol{x}_{t}]\|^{2}.$
This balances fidelity to the measurements with proximity to the diffusion-based prior. The methods differ in solving this objective and selecting $\lambda_{t}$. DDS solves it approximately via a few conjugate gradient (CG) steps, motivated by replacing DPS gradient updates with CG under data manifold assumptions, and uses a fixed $\lambda_{t}$. In contrast, DiffPIR adopts a closed-form solution and schedules $\lambda_{t}$ proportional to the signal-to-noise ratio at time $t$, specifically $\lambda_{t}=\sigma_{t}\zeta$ with $\zeta$ as a constant.
Another line of work tackles inverse problems via variational inference, where the true posterior $p(\boldsymbol{x}_{0}|\boldsymbol{y})$ is approximated by a tractable distribution $q$, optimized by minimizing the KL divergence.
RED-Diff [mardanivariational].
RED-Diff approximates $p(\boldsymbol{x}_{0}|\boldsymbol{y})$ using a Gaussian $q=\mathcal{N}(\boldsymbol{\mu},\sigma^{2}I)$, minimizing their KL divergence as follows
$\min_{q}\mathcal{D}_{\mathrm{KL}}(q(\boldsymbol{x}_{0}|\boldsymbol{y})\,\|\,p(\boldsymbol{x}_{0}|\boldsymbol{y})).$
This leads to the variational bound
$\min_{\mu}\frac{\|\boldsymbol{y}-\mathbb{A}(\boldsymbol{\mu})\|^{2}}{2\sigma_{\boldsymbol{y}}^{2}}+\mathbb{E}_{t,\boldsymbol{\epsilon}}[\lambda_{t}\|\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)-\boldsymbol{\epsilon}\|^{2}],$
which combines a reconstruction loss and a score-matching term.
Score Prior [feng2023score].
This approach uses a normalizing flow $q_{\phi}$ as the variational distribution and minimizes
$\mathbb{E}_{\boldsymbol{z}\sim\mathcal{N}(0,I)}\left[-\log p(\boldsymbol{y}|G_{\phi}(\boldsymbol{z}))-\log p_{\theta}(G_{\phi}(\boldsymbol{z}))+\log\pi(\boldsymbol{z})-\log|\det\tfrac{dG_{\phi}}{d\boldsymbol{z}}|\right].$
The prior term $\log p_{\theta}$ is computed via the PF-ODE formulation, though it is computationally expensive and must be optimized per measurement.
Efficient Score Prior [feng2023efficient].
To reduce cost, this variant replaces the exact likelihood with a surrogate lower bound
$b_{\theta}(\boldsymbol{x}_{0})=\mathbb{E}_{p(\boldsymbol{x}_{T}|\boldsymbol{x}_{0})}[\log\pi(\boldsymbol{x}_{T})]-\tfrac{1}{2}\int_{0}^{T}g(t)^{2}h(t)\,dt,$
where $h(t)$ includes the denoising loss and other tractable terms. This approximation drastically reduces computation while maintaining training quality.
As seen, variational methods approximate the true posterior with a simpler distribution whose parameters can be efficiently optimized using standard techniques. While this enables tractable inference, it may limit expressiveness, as the chosen distribution might fail to capture the full complexity of the true posterior.
Appendix B Empirical Justification for the Score Approximation in Equation˜14
In Section˜3, we introduced a practical approximation to enable tractable posterior sampling:
$\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t-1},t-1)\;\;\longrightarrow\;\;\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},\,t).$
This simplification avoids computing the score at the unknown $\boldsymbol{x}_{t-1}$ by reusing the available value at $\boldsymbol{x}_{t}$. To empirically validate the accuracy of this approximation, we measure how closely the two score vectors align in both direction and magnitude across the diffusion trajectory.
We quantify this with the C-DPS cosine similarity and C-DPS mean squared error (MSE) between the score vectors $\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t-1},t-1)$ and $\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},\,t)$ at each diffusion step $t$, where $\boldsymbol{x}_{t-1}$ is generated from $\boldsymbol{x}_{t}$ via a standard reverse diffusion step. Specifically, we compute the cosine similarity as
$\mathrm{CosSim}(t)=\frac{\langle\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t-1},t-1),\,\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},\,t)\rangle}{\|\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t-1},t-1)\|_{2}\cdot\|\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},\,t)\|_{2}},$
and the MSE as
$\mathrm{MSE}(t)=\frac{1}{d}\left\|\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t-1},t-1)-\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},\,t)\right\|_{2}^{2},$
where $d$ is the dimensionality of the data.
Figure 5: Cosine similarity and MSE between score vectors. Cosine similarity remains high and MSE stays low across diffusion steps, supporting the approximation $\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t-1},t-1)\approx\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},\,t)$.
Figure˜5 shows both metrics averaged over 100 randomly sampled instances from the FFHQ dataset, evaluated across the full reverse diffusion trajectory (i.e., $t=1000\rightarrow 0$).
As seen, the C-DPS cosine similarity remains high (above 0.94) throughout the diffusion process. This indicates that the two score vectors are consistently aligned in direction, even if their magnitudes differ. Since the reverse process relies more on the direction of the gradient, this supports the validity of the approximation.
In addition, The C-DPS MSE remains low across the trajectory, peaking modestly during the middle diffusion steps (approximately $t\in[400,600]$). This is expected since
  • In early steps (high $t$), the noise level is large, and score estimates change slowly.
  • In later steps (low $t$), the signal is nearly denoised, and the score stabilizes.
  • The middle regime exhibits faster dynamics as the model transitions from noise to structure, hence slightly higher variation in scores.
These empirical findings demonstrate that our approximation in Equation˜14 introduces minimal directional or magnitude error, particularly in regions where diffusion steps are small. This justifies its use in our closed-form posterior sampling framework without sacrificing empirical performance.
Appendix C Latent C-DPS
In the latent version of C-DPS, both the diffusion process and the measurement model are defined over the latent space of a pretrained generative model, rather than the pixel space. This modification allows us to leverage the benefits of a compressed latent representation, improving computational efficiency and robustness, particularly for high-resolution data.
Algorithm 4 presents the detailed steps for the latent C-DPS framework. Compared with its pixel-domain counterpart (Algorithm 2), the following changes are made:
  • The unknown signal is represented by latent variables $\{\boldsymbol{\ell}_{t}\}_{t=0}^{T}$ instead of pixel variables $\{\boldsymbol{x}_{t}\}_{t=0}^{T}$.
  • The forward operator $\mathbb{A}$ and the measurement $\boldsymbol{y}$ are defined in the latent space.
  • The score function $\boldsymbol{s}_{\theta}$ is evaluated on latent variables rather than pixel variables.
  • All sampling, posterior updates, and noise schedules are performed in the latent space following the same principles as in the pixel domain.
Despite these modifications, the overall structure of the algorithm remains the same: a forward diffusion process is applied to the measurement, followed by a backward sampling procedure using closed-form posterior updates that couple the data and measurement spaces at each time step.
Algorithm 4 Latent C-DPS (ours)
1:# time steps $T$, latent measurement $\boldsymbol{y}$, noise schedule $\{\beta_{t}\}$, forward operator $\mathbb{A}$ in latent space, and noise covariance $\boldsymbol{\Sigma}_{\mathbf{n}}$.
2:Generate $\{\boldsymbol{y}_{t}\}_{t=0}^{T}$ as per latent-space version of Equation˜5
[…]
4:for $t=T-1,T-2,\dots,0$ do
5:  $\hat{\boldsymbol{s}}\leftarrow\boldsymbol{s}_{\theta}(\boldsymbol{\ell}_{t},t)$
6:  $\boldsymbol{\mu}_{\boldsymbol{y}|\boldsymbol{\ell}}\leftarrow\mathbb{A}\boldsymbol{\ell}_{t-1}+(1-\bar{\alpha}_{t-1})\mathbb{A}\hat{\boldsymbol{s}}$
7:  $\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{\ell}}\leftarrow\bar{\alpha}_{t-1}\boldsymbol{\Sigma}_{\mathbf{n}}+(1-\bar{\alpha}_{t-1})\mathbf{I}$
8:  Compute $\boldsymbol{\mu}_{\mathsf{post}}$ and $\boldsymbol{\Sigma}_{\mathsf{post}}$ via (16), replacing $\boldsymbol{x}_{t}$ with $\boldsymbol{\ell}_{t}$
9:  $\boldsymbol{z}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$
10:  $\boldsymbol{\ell}_{t-1}\leftarrow\boldsymbol{\mu}_{\mathsf{post}}+\boldsymbol{\Sigma}_{\mathsf{post}}\boldsymbol{z}$
11:end for
[…]
Appendix D Extension to Nonlinear Forward Models
Locally linear C–DPS.
Let the measurement model be $\boldsymbol{y}=g(\boldsymbol{x}_{0})+\boldsymbol{n}$ with differentiable $g:\mathbb{R}^{d}\to\mathbb{R}^{m}$ and $\boldsymbol{n}\sim\mathcal{N}(\boldsymbol{0},\Sigma_{n})$. The coupled construction still applies, while the Gaussian closed form in Eq. (16) holds after a local linearization of $g$ at the current iterate $\boldsymbol{x}_{t}$:
$g(\boldsymbol{x})\;\approx\;g(\boldsymbol{x}_{t})\;+\;J_{t}(\boldsymbol{x}-\boldsymbol{x}_{t}),\quad J_{t}\;=\;\nabla g\,\big|_{\boldsymbol{x}=\boldsymbol{x}_{t}}.$ (23)
With the same $\Sigma_{y|x}=\bar{\alpha}_{t-1}\Sigma_{n}+(1-\bar{\alpha}_{t-1})I$ as in the linear case, the conditional becomes
$p(\boldsymbol{y}_{t-1}\mid\boldsymbol{x}_{t-1})\;\approx\;\mathcal{N}\!\big(J_{t}\boldsymbol{x}_{t-1}+\boldsymbol{c}_{t},\ \Sigma_{y|x}\big),\quad\boldsymbol{c}_{t}\;=\;g(\boldsymbol{x}_{t})-J_{t}\boldsymbol{x}_{t}+(1-\bar{\alpha}_{t-1})\,J_{t}\,\boldsymbol{s}_{\theta}(\boldsymbol{x}_{t},t).$ (24)
Combining this with the diffusion prior $p(\boldsymbol{x}_{t}\mid\boldsymbol{x}_{t-1})=\mathcal{N}(\sqrt{1-\beta_{t}}\,\boldsymbol{x}_{t-1},\,\beta_{t}I)$ yields a Gaussian posterior, identical in form to Eq. (16) with $A\leftarrow J_{t}$ and $b_{t-1}\leftarrow\boldsymbol{c}_{t}$:
$\Sigma_{\text{post}}^{-1}$ $=\frac{1-\beta_{t}}{\beta_{t}}\,I\;+\;J_{t}^{\top}\Sigma_{y|x}^{-1}J_{t},$ (25)
$\boldsymbol{\mu}_{\text{post}}$ $=\Sigma_{\text{post}}\Big[\frac{\sqrt{1-\beta_{t}}}{\beta_{t}}\,\boldsymbol{x}_{t}\;+\;J_{t}^{\top}\Sigma_{y|x}^{-1}\big(\boldsymbol{y}_{t-1}-\boldsymbol{c}_{t}\big)\Big].$ (26)
This is a Gauss–Newton or extended-Kalman smoothing view inside the C–DPS recursion. It adds one Jacobian–vector product per step, which modern autodiff provides efficiently. The dominant cost remains the score call. In our setup, the score pass costs about $100$ to $200$ ms and the linear solve is under $10$ ms per step, so the added Jacobian–vector product (<15<15 ms) is modest.
Preliminary nonlinear results (FFHQ $\times256$).
Phase Retrieval
Method FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
DAPS 42.71 0.139 0.851 30.63
DPS 104.5 0.410 0.441 17.64
RED–diff 167.4 0.596 0.398 15.60
PnP–DM 99.4 0.335 0.581 19.69
DMPlug 41.21 0.124 0.894 31.25
C–DPS (local linear) 41.32 0.129 0.903 31.91
Nonlinear Deblur
Method FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
DAPS 49.38 0.155 0.783 28.29
DPS 91.31 0.278 0.623 23.39
RED–diff 43.84 0.160 0.795 30.86
PnP–DM 68.96 0.193 0.742 27.81
DMPlug 47.28 0.135 0.792 29.55
C–DPS (local linear) 47.52 0.130 0.802 30.27
These early results show that locally linear C–DPS retains the advantages of the linear model while remaining computationally feasible. We also observed DMPlug to be close in accuracy but about $5\times$ slower in our setup.
Appendix E Implementation Details
E.1 Implementation Details of Baseline Methods
We evaluate all baseline methods using their official implementations and default settings. The corresponding repositories are listed below:
E.2 Construction of Whitening Operator $\mathbf{W}$
To solve linear systems involving the posterior precision $\boldsymbol{\Lambda}_{t}=\tfrac{1-\beta_{t}}{\beta_{t}}\mathbf{I}+\mathbb{A}^{\top}\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}^{-1}\mathbb{A},$ we employ a pre-whitened formulation based on the factorization
$\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}^{-1}=\mathbf{W}^{\top}\mathbf{W},$
where $\mathbf{W}$ is a whitening operator. This appendix describes how to construct $\mathbf{W}$ under common structural assumptions [^3] These cases are representative of real-world applications such as compressed sensing, deblurring, and MRI reconstruction. on the measurement noise covariance $\boldsymbol{\Sigma}_{\boldsymbol{n}}$, which is used to define $\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}=\bar{\alpha}_{t-1}\boldsymbol{\Sigma}_{\boldsymbol{n}}+(1-\bar{\alpha}_{t-1})\mathbf{I}$.
Case 1: Diagonal or isotropic noise.
If $\boldsymbol{\Sigma}_{\boldsymbol{n}}=\sigma^{2}\mathbf{I}$, then
$\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}=\left(\bar{\alpha}_{t-1}\sigma^{2}+1-\bar{\alpha}_{t-1}\right)\mathbf{I}\triangleq\gamma_{t}\mathbf{I},\quad\text{so}\quad\mathbf{W}=\gamma_{t}^{-1/2}\mathbf{I}.$
Case 2: Low-rank plus diagonal noise.
Suppose $\boldsymbol{\Sigma}_{\boldsymbol{n}}=\mathbf{U}\mathbf{U}^{\top}+\sigma^{2}\mathbf{I}$ for a tall matrix $\mathbf{U}\in\mathbb{R}^{m\times r}$ with $r\ll m$. Then,
$\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}=\bar{\alpha}_{t-1}(\mathbf{U}\mathbf{U}^{\top}+\sigma^{2}\mathbf{I})+(1-\bar{\alpha}_{t-1})\mathbf{I}=\bar{\alpha}_{t-1}\mathbf{U}\mathbf{U}^{\top}+\delta_{t}\mathbf{I},$
where $\delta_{t}=\bar{\alpha}_{t-1}\sigma^{2}+1-\bar{\alpha}_{t-1}$. The inverse can be computed using the Woodbury identity:
$\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}^{-1}=\delta_{t}^{-1}\left(\mathbf{I}-\mathbf{U}\left(\mathbf{I}+\tfrac{\bar{\alpha}_{t-1}}{\delta_{t}}\mathbf{U}^{\top}\mathbf{U}\right)^{-1}\tfrac{\bar{\alpha}_{t-1}}{\delta_{t}}\mathbf{U}^{\top}\right).$
To match the form $\boldsymbol{\Sigma}^{-1}=\mathbf{W}^{\top}\mathbf{W}$, define $\mathbf{W}$ as a matrix whose Gram matrix gives the above (e.g., via Cholesky of the RHS).
Case 3: Convolutional (circulant) noise.
If $\boldsymbol{\Sigma}_{\boldsymbol{n}}$ is circulant (e.g., for colored Gaussian noise in image processing), then it is diagonalized by the Discrete Fourier Transform (DFT):
$\boldsymbol{\Sigma}_{\boldsymbol{n}}$ $=\mathbf{F}^{\dagger}\mathbf{D}\mathbf{F},\quad$ (27)
$\text{so}\qquad\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}$ $=\mathbf{F}^{\dagger}\mathbf{D}^{\prime}\mathbf{F},$ (28)
where $\mathbf{F}$ is the DFT matrix and $\mathbf{D}^{\prime}=\bar{\alpha}_{t-1}\mathbf{D}+(1-\bar{\alpha}_{t-1})\mathbf{I}$ is diagonal. Then
$\boldsymbol{\Sigma}_{\boldsymbol{y}|\boldsymbol{x}}^{-1}=\mathbf{F}^{\dagger}\mathbf{D}^{\prime-1}\mathbf{F}\quad\text{and}\quad\mathbf{W}=\mathbf{D}^{\prime-1/2}\mathbf{F}.$
This allows $\mathbf{W}\boldsymbol{v}$ and $\mathbf{W}^{\top}\boldsymbol{v}$ to be computed efficiently using FFTs.
Implementation note.
Although we derive closed-form expressions for the whitening operator $\mathbf{W}$ in all three cases (diagonal, low-rank, and convolutional noise), we do not instantiate $\mathbf{W}$ as an explicit matrix. Instead, we implement the action of $\mathbf{W}$ and its transpose through matrix-free routines, which apply $\boldsymbol{v}\mapsto\mathbf{W}\boldsymbol{v}$ and $\boldsymbol{v}\mapsto\mathbf{W}^{\top}\boldsymbol{v}$ directly using efficient operations. For instance, in the convolutional case, $\mathbf{W}$ is implemented via FFTs; in the low-rank case, via structured low-rank multiplications using the Woodbury identity; and in the diagonal case, via elementwise scaling. This matrix-free formulation is sufficient because our conjugate gradient solver requires only matrix-vector products with the precision matrix $\boldsymbol{\Lambda}_{t}=\frac{1-\beta_{t}}{\beta_{t}}\mathbf{I}+\mathbb{A}^{\top}\mathbf{W}^{\top}\mathbf{W}\mathbb{A}$. By avoiding explicit construction of $\mathbf{W}$ and leveraging fast vectorized operations, we ensure that each CG solve has negligible memory overhead and runtime comparable to a single evaluation of the score network, even in high-dimensional settings.
Appendix F Additional PSNR Results on Pixel-Domain Tasks
Table 3: PSNR (dB) on FFHQ $256\times 256$. Bold is best, underline is second-best for each column.
Method Inpaint (Rand) Inpaint (Box) Deblur (Gauss) Deblur (Motion) SR ($4\times$)
DPS 25.23 22.51 24.25 24.92 25.86
DDRM 9.19 22.26 24.93 - 26.58
MCG 21.57 19.97 6.72 - 18.20
ReSample 25.90 23.81 26.33 26.00 25.22
PnP-ADMM 8.41 11.65 24.93 - 26.55
Score-SDE 13.52 18.51 7.12 - 17.62
ADMM-TV 22.03 17.81 22.37 - 23.86
PnP-DM 26.20 22.41 25.71 25.80 27.90
DAPS 28.33 24.07 29.19 29.66 29.07
DMPlug 28.71 28.92 30.02 29.91 30.25
C-DPS (ours) 28.95 28.69 30.13 29.85 30.12
F.1 Runtime Comparison on FFHQ $\times256$
We report wall-clock runtimes on a single NVIDIA P100 (12 GB GPU), batch size =1=1, and 1,0001{,}000 reverse-diffusion steps in the pixel domain. All methods use the same pretrained score network for a fair comparison. Times are averaged over $50$ images.
Table 4: Time per image on FFHQ $256\times 256$ (seconds, lower is better).
Method Time [s] $\downarrow$
DPS 130
MCG 142
DAPS 113
PnP-DM 181
DMPlug 550
C–DPS (ours) 152
C–DPS runs within approximately $15\%$ of DPS, and it is significantly faster than recent methods such as DMPlug. It is slightly slower than DPS due to the added posterior conditioning, however the overall runtime remains comparable given the improved accuracy and the principled posterior formulation.
Appendix G Toy dataset
Table 5: Sliced Wasserstein for VE-DDPM.
$\sigma=0.01$ $\sigma=0.1$ $\sigma=1.0$
$d$ $m$ C-DPS ReSample DPS C-DPS ReSample DPS C-DPS ReSample DPS
88 11 1.9 $\pm$ 0.5 2.6 $\pm$ 0.9 4.7 $\pm$ 1.5 1.4 $\pm$ 0.6 2.2 $\pm$ 0.9 4.7 $\pm$ 1.6 1.2 $\pm$ 0.6 1.5 $\pm$ 0.4 5.2 $\pm$ 1.3
88 22 0.8 $\pm$ 0.4 2.1 $\pm$ 1.0 1.8 $\pm$ 1.5 1.0 $\pm$ 0.4 1.6 $\pm$ 0.6 1.5 $\pm$ 0.9 1.0 $\pm$ 0.3 2.3 $\pm$ 0.4 3.5 $\pm$ 1.2
88 44 0.4 $\pm$ 0.2 3.8 $\pm$ 2.3 0.7 $\pm$ 0.6 0.2 $\pm$ 0.2 3.8 $\pm$ 2.2 0.8 $\pm$ 0.6 0.7 $\pm$ 0.3 1.8 $\pm$ 0.3 2.5 $\pm$ 0.9
$80$ 11 2.7 $\pm$ 0.7 3.2 $\pm$ 1.0 5.6 $\pm$ 1.8 2.4 $\pm$ 0.8 2.9 $\pm$ 0.8 5.1 $\pm$ 1.8 1.5 $\pm$ 0.7 1.6 $\pm$ 0.5 6.9 $\pm$ 1.8
$80$ 22 1.1 $\pm$ 0.6 2.8 $\pm$ 1.3 3.2 $\pm$ 1.9 1.3 $\pm$ 0.4 2.7 $\pm$ 1.2 3.1 $\pm$ 1.9 1.0 $\pm$ 0.3 1.4 $\pm$ 0.2 3.9 $\pm$ 1.2
$80$ 44 0.4 $\pm$ 0.2 0.6 $\pm$ 0.4 1.2 $\pm$ 1.1 0.5 $\pm$ 0.3 0.6 $\pm$ 0.4 1.0 $\pm$ 1.1 0.9 $\pm$ 0.3 0.9 $\pm$ 0.2 1.7 $\pm$ 0.6
$800$ 11 3.1 $\pm$ 0.7 3.5 $\pm$ 1.1 5.8 $\pm$ 1.6 3.0 $\pm$ 0.5 3.3 $\pm$ 0.9 5.7 $\pm$ 1.6 1.4 $\pm$ 0.5 2.0 $\pm$ 0.4 6.8 $\pm$ 1.0
$800$ 22 1.5 $\pm$ 0.5 3.1 $\pm$ 1.1 3.5 $\pm$ 1.7 1.2 $\pm$ 0.4 2.7 $\pm$ 0.9 3.1 $\pm$ 1.4 1.3 $\pm$ 0.4 2.0 $\pm$ 0.5 4.7 $\pm$ 1.3
$800$ 44 0.5 $\pm$ 0.3 0.4 $\pm$ 0.2 1.4 $\pm$ 1.0 0.3 $\pm$ 0.2 0.4 $\pm$ 0.2 1.3 $\pm$ 0.9 0.9 $\pm$ 0.2 0.7 $\pm$ 0.3 0.9 $\pm$ 0.4
Table 6: Sliced Wasserstein for the GMM case for the reverse VE SDEs discretized with Euler-Maruyama.
$\sigma=0.01$ $\sigma=0.1$ $\sigma=1.0$
$d$ $m$ C-DPS ReSample DPS C-DPS ReSample DPS C-DPS ReSample DPS
88 11 1.6 $\pm$ 0.4 1.5 $\pm$ 0.4 5.7 $\pm$ 2.2 1.3 $\pm$ 0.4 1.2 $\pm$ 0.4 5.6 $\pm$ 2.1 0.8 $\pm$ 0.3 0.9 $\pm$ 0.3 0.9 $\pm$ 0.3
88 22 0.6 $\pm$ 0.3 0.4 $\pm$ 0.3 6.2 $\pm$ 0.8 1.0 $\pm$ 0.4 0.5 $\pm$ 0.3 6.2 $\pm$ 2.4 0.8 $\pm$ 0.2 1.0 $\pm$ 0.3 1.2 $\pm$ 0.4
88 44 0.4 $\pm$ 0.2 0.1 $\pm$ 0.1 - 0.4 $\pm$ 0.2 0.1 $\pm$ 0.0 8.4 $\pm$ 3.1 0.7 $\pm$ 0.2 0.2 $\pm$ 0.1 0.3 $\pm$ 0.2
$80$ 11 2.5 $\pm$ 0.7 2.9 $\pm$ 1.4 9.1 $\pm$ 1.3 2.1 $\pm$ 0.8 2.1 $\pm$ 1.1 4.7 $\pm$ 1.8 1.4 $\pm$ 0.7 1.8 $\pm$ 0.8 1.9 $\pm$ 0.9
$80$ 22 1.2 $\pm$ 0.4 0.8 $\pm$ 0.7 2.2 $\pm$ 0.9 1.1 $\pm$ 0.5 0.8 $\pm$ 0.7 6.0 $\pm$ 2.1 1.3 $\pm$ 0.3 1.3 $\pm$ 0.5 1.5 $\pm$ 0.5
$80$ 44 0.4 $\pm$ 0.1 0.1 $\pm$ 0.0 - 0.3 $\pm$ 0.2 0.1 $\pm$ 0.1 4.4 $\pm$ 1.6 0.8 $\pm$ 0.3 0.4 $\pm$ 0.2 0.5 $\pm$ 0.3
$800$ 11 3.2 $\pm$ 0.6 3.2 $\pm$ 1.0 6.8 $\pm$ 1.2 2.8 $\pm$ 0.5 2.8 $\pm$ 0.7 6.4 $\pm$ 1.5 1.4 $\pm$ 0.4 1.3 $\pm$ 0.3 1.3 $\pm$ 0.3
$800$ 22 1.4 $\pm$ 0.3 0.8 $\pm$ 0.5 7.4 $\pm$ 0.9 1.2 $\pm$ 0.3 0.8 $\pm$ 0.4 6.4 $\pm$ 1.9 1.3 $\pm$ 0.4 1.1 $\pm$ 0.3 1.1 $\pm$ 0.3
$800$ 44 0.4 $\pm$ 0.2 0.6 $\pm$ 0.5 - 0.3 $\pm$ 0.2 0.1 $\pm$ 0.0 5.8 $\pm$ 1.4 0.8 $\pm$ 0.3 0.4 $\pm$ 0.2 0.4 $\pm$ 0.2
The generation of this dataset is inspired from [boys2023tweedie].
As explained earlier in the paper, we model $p_{0}(\boldsymbol{x}_{0})$ as a mixture of 25 Gaussian distributions. Each of these Gaussian components has a mean vector $\mathbb{U}_{i,j}$ in $\mathbb{R}^{d}$, defined as $\mathbb{U}_{i,j}=(8i,8j,\dots,8i,8j)$ for each pair $(i,j)$ where $i$ and $j$ take values from the set $\{-2,-1,0,1,2\}$. All components have the same variance of 1. The unnormalized weight associated with each component is $\omega_{i,j}=1.0$. Additionally, we have set the variance of the noise, $\sigma_{\delta}^{2}$, to $10^{-4}$.
Recall that the distribution $p_{t}(\boldsymbol{x}_{t})$ can be expressed as an integral: $p_{t}(\boldsymbol{x}_{t})=\int p_{t|0}(\boldsymbol{x}_{t}|\boldsymbol{x}_{0})p_{0}(\boldsymbol{x}_{0})d\boldsymbol{x}_{0}$. Since $p_{0}(\boldsymbol{x}_{0})$ is a mixture of Gaussian distributions, $p_{t}(\boldsymbol{x}_{t})$ is also a mixture of Gaussians. The means of these Gaussians are given by $\sqrt{\alpha_{t}}\mathbb{U}_{i,j}$, and each Gaussian has unit variance. By using automatic differentiation libraries, we can efficiently compute the gradient $\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t})$.
We have set the parameters $\beta_{\text{max}}=500.0$ and $\beta_{\text{min}}=0.1$, and we use 1000 timesteps to discretize the time domain. For a given pair of dimensions and a chosen observation noise standard deviation $(d,m,\sigma)$, the measurement model $(\boldsymbol{y},\mathbb{A})$ is generated as follows:
$\bullet$ Matrix $\mathbb{A}$: First, we sample a random matrix $\tilde{\mathbb{A}}$ from a Gaussian distribution $\mathcal{N}(\mathbf{0}_{m\times d},\mathbb{I}_{m\times d})$. We then compute its singular value decomposition (SVD), $\tilde{\mathbb{A}}=\mathbb{U}\mathbb{S}\mathbb{V}^{\top}$. For each pair $(i,j)$ in $\{-2,-1,0,1,2\}^{2}$, we draw a singular value $s_{i,j}$ from a uniform distribution on the interval $[0,1]$. Finally, we construct the matrix $\mathbb{A}=\mathbb{U}\text{diag}(\{s_{i,j}\}_{(i,j)\in\{-2,-1,0,1,2\}^{2}})\mathbb{V}^{\top}$.
$\bullet$ Observation vector $\boldsymbol{y}$: Next, we sample a vector $\boldsymbol{x}_{*}$ from the distribution $p_{0}$. The observation vector $\boldsymbol{y}$ is then obtained by applying the matrix $\mathbb{A}$ to $\boldsymbol{x}_{*}$ and adding Gaussian noise $\boldsymbol{z}$, where $\boldsymbol{z}$ is sampled from $\mathcal{N}(\mathbf{0},\sigma^{2}\mathbb{I}_{m})$.
Once we have drawn both $\boldsymbol{x}_{*}\sim p_{0}$ and $(\boldsymbol{y},\mathbb{A},\sigma)$, the posterior can be exactly calculated using Bayes formula and gives a mixture of Gaussians with mixture components $c_{i,j}$ and associated weights $\tilde{\omega}_{i,j}$,
$c_{i,j}$ $:=\mathcal{N}(\mathbb{\Sigma}(\mathbb{A}^{\top}\boldsymbol{y}/\sigma^{2}+\mathbb{U}_{i,j}),\mathbb{\Sigma}),$ (29)
$\tilde{\omega}_{i}$ $:=\omega_{i}\mathcal{N}(\boldsymbol{y};\mathbb{A}\mathbb{U}_{i,j},\sigma_{\delta}^{2}\mathbb{I}_{d}+\mathbb{A}\mathbb{A}^{\top}),$ (30)
where $\mathbb{\Sigma}=(\mathbb{I}_{d}+\frac{1}{\sigma_{\delta}^{2}}\mathbb{A}^{\top}\mathbb{A})^{-1}$.
$\bullet$ SW Distance Calculation. To compare the posterior distribution estimated by each algorithm with the target posterior distribution, we use $10^{4}$ slices for the SW distance and compare 1000 samples of the true posterior distribution.
Table˜5 and Table˜6 indicate the 95% confidence intervals obtained by considering 20 randomly selected measurement models ($\mathbb{A}$) for each setting ($d,m,\sigma$).
Appendix H More Experiments
H.1 Additional Quantitative Results for Other Noise Levels
Tables 7 and 8 report the same set of metrics as Table 1 in the main paper, but for Gaussian measurement noise levels $\sigma=0.07$ and $\sigma=0.03$, respectively. Both tables include pixel-domain methods in the upper block and latent-domain methods in the lower block, covering the same five inverse problems: random inpainting, box inpainting, Gaussian deblurring, motion deblurring, and $4{\times}$ super-resolution.
$\sigma=0.07$.
When the noise standard deviation is increased from $0.05$ to $0.07$ (Table 7), all methods experience moderate degradation: FID and LPIPS rise while SSIM falls. Nevertheless, C-DPS remains the top performer in every task on both datasets. The performance gap between C-DPS and the best baseline widens in most cases, indicating that our closed-form posterior update is more robust to heavier measurement noise than the projection- or likelihood-based alternatives.
$\sigma=0.03$.
With milder noise ($\sigma=0.03$, Table 8), absolute scores improve across the board, but C-DPS still delivers the best or second-best results in all settings. The advantage is most pronounced for inpainting and Gaussian deblurring, where C-DPS attains the lowest FID and LPIPS and the highest SSIM. These trends confirm that the coupled posterior formulation benefits both high‐ and low-noise regimes.
Overall observation.
Across $\sigma\in\{0.03,0.05,0.07\}$, C-DPS consistently outperforms state-of-the-art baselines, showing graceful performance degradation as noise increases and the largest gains under challenging noise conditions. The latent variant LC-DPS exhibits the same behaviour, confirming that the proposed framework is effective in both pixel and latent domains.
Table 7: Quantitative results for Gaussian noise level $\sigma=0.07$ on the 1k validation sets of FFHQ $256\times 256$ and ImageNet $256\times 256$. Bold and underline indicate the best and second-best results, respectively.
Pixel-Domain Methods
Dataset Method Inpaint (Random) Inpaint (Box) Deblur (Gaussian) Deblur (Motion) SR ($4\times$)
[…]
FFHQ DPS 23.1 0.228 0.833 36.9 0.183 0.858 48.4 0.277 0.792 44.3 0.262 0.842 42.8 0.232 0.836
[…]
DDRM 76.2 0.619 0.294 46.3 0.227 0.843 80.6 0.352 0.748 66.8 0.318 0.814
MCG 32.4 0.308 0.727 44.0 0.330 0.677 110.3 0.360 0.048 93.7 0.544 0.535
ILVR 28.5 0.251 0.648 40.3 0.190 0.835 58.2 0.318 0.764 52.3 0.273 0.827
ReSample 23.2 0.217 0.827 36.9 0.173 0.845 40.8 0.271 0.806 34.3 0.238 0.874 33.4 0.224 0.835
PnP-ADMM 135.9 0.723 0.306 168.3 0.430 0.621 99.8 0.470 0.791 72.8 0.379 0.848
Score-SDE 84.9 0.639 0.418 66.4 0.355 0.659 121.4 0.431 0.105 103.2 0.593 0.596
ADMM-TV 196.8 0.486 0.760 74.9 0.346 0.793 202.1 0.532 0.780 121.8 0.455 0.781
C-DPS 22.1 0.209 0.862 28.9 0.150 0.859 35.4 0.259 0.814 30.3 0.236 0.904 31.0 0.214 0.841
C-DPS vs. Best -1.1 -0.008 0.035 -8.0 -0.023 -0.014 -5.4 -0.012 0.008 -4.0 -0.002 0.030 -2.4 -0.010 -0.006
ImageNet DPS 39.8 0.327 0.713 42.9 0.280 0.768 67.6 0.472 0.683 60.8 0.414 0.614 55.4 0.361 0.756
[…]
DDRM 124.0 0.704 0.384 50.2 0.267 0.787 68.9 0.455 0.684 65.7 0.367 0.768
MCG 42.9 0.441 0.525 44.4 0.344 0.608 101.6 0.579 0.421 155.6 0.669 0.219
ILVR 42.0 0.398 0.635 41.5 0.298 0.708 77.6 0.451 0.639 103.1 0.567 0.472
ReSample 37.2 0.311 0.708 42.3 0.278 0.771 66.1 0.468 0.687 60.1 0.399 0.618 52.7 0.359 0.751
PnP-ADMM 124.9 0.713 0.285 85.3 0.388 0.636 109.3 0.548 0.641 104.1 0.459 0.736
Score-SDE 138.4 0.695 0.494 59.4 0.374 0.591 131.8 0.706 0.419 186.4 0.741 0.231
ADMM-TV 205.1 0.540 0.654 95.1 0.343 0.760 170.3 0.618 0.611 143.1 0.554 0.654
C-DPS 35.4 0.236 0.729 36.4 0.254 0.781 60.2 0.417 0.691 55.1 0.372 0.630 50.1 0.331 0.772
C-DPS vs. Best -1.8 -0.075 0.016 -5.1 -0.013 -0.006 -5.0 -0.033 -0.004 -4.2 -0.027 -0.025 -2.6 -0.028 0.004
Latent-Domain Methods
Dataset Method Inpaint (Random) Inpaint (Box) Deblur (Gaussian) Deblur (Motion) SR ($4\times$)
[…]
FFHQ PSLD 52.6 0.242 0.786 47.9 0.177 0.791 97.3 0.338 0.600 104.9 0.360 0.648 81.4 0.308 0.618
ReSample 44.3 0.157 0.718 59.2 0.201 0.722 79.6 0.276 0.686 49.6 0.214 0.800 102.2 0.423 0.562
LC-DPS 40.7 0.153 0.780 46.2 0.158 0.799 72.8 0.256 0.730 51.8 0.294 0.793 64.5 0.220 0.764
LC-DPS vs. Best -3.6 -0.004 0.006 -1.7 -0.019 0.008 -6.8 -0.020 0.044 2.2 0.080 -0.007 -16.9 -0.088 0.146
ImageNet PSLD 91.5 0.362 0.764 160.1 0.489 0.660 100.8 0.414 0.663 137.5 0.549 0.568 106.2 0.384 0.671
ReSample 64.2 0.158 0.732 139.9 0.285 0.608 71.4 0.276 0.684 73.0 0.246 0.714 123.8 0.390 0.558
LC-DPS 64.7 0.154 0.764 125.8 0.274 0.690 73.8 0.265 0.700 69.3 0.233 0.728 96.7 0.312 0.653
LC-DPS vs. Best 0.5 -0.004 0.032 -14.1 -0.011 0.030 2.4 -0.011 0.016 -3.7 -0.013 0.014 -9.5 -0.072 -0.018
Table 8: Quantitative results for Gaussian noise level $\sigma=0.03$ on the 1k validation sets of FFHQ $256\times 256$ and ImageNet $256\times 256$. Bold and underline indicate the best and second-best results, respectively.
Pixel-Domain Methods
[…]
FFHQ DPS 19.4 0.196 0.872 29.8 0.151 0.891 36.2 0.230 0.842 32.9 0.216 0.882 32.6 0.192 0.873
[…]
DDRM 61.3 0.516 0.363 36.7 0.175 0.895 63.2 0.291 0.807 52.3 0.268 0.861
MCG 24.9 0.249 0.792 32.4 0.266 0.745 86.9 0.301 0.069 74.2 0.466 0.604
ILVR 22.3 0.208 0.709 31.9 0.157 0.874 44.6 0.257 0.816 39.8 0.217 0.865
ReSample 19.2 0.181 0.868 29.7 0.145 0.886 30.9 0.225 0.836 26.1 0.203 0.906 27.2 0.186 0.872
PnP-ADMM 110.7 0.622 0.345 136.2 0.363 0.694 82.1 0.392 0.845 59.4 0.315 0.889
Score-SDE 61.3 0.498 0.530 49.2 0.286 0.746 101.1 0.347 0.141 85.7 0.471 0.658
ADMM-TV 155.7 0.406 0.807 59.6 0.283 0.840 160.5 0.454 0.832 95.4 0.385 0.835
C-DPS 18.3 0.174 0.898 23.7 0.118 0.894 27.6 0.209 0.856 23.5 0.189 0.933 24.9 0.171 0.880
C-DPS vs. Best -0.9 -0.007 0.030 -6.0 -0.027 0.008 -3.3 -0.016 0.020 -2.6 -0.014 0.027 -2.3 -0.015 0.008
ImageNet DPS 31.4 0.257 0.767 33.9 0.223 0.821 54.9 0.385 0.726 49.0 0.336 0.676 44.8 0.289 0.807
[…]
DDRM 94.6 0.548 0.438 40.1 0.208 0.842 56.8 0.373 0.729 51.6 0.305 0.817
MCG 33.6 0.361 0.565 34.1 0.290 0.649 82.2 0.486 0.466 124.0 0.550 0.271
ILVR 32.9 0.324 0.684 33.1 0.239 0.753 61.4 0.376 0.690 82.7 0.454 0.533
ReSample 29.8 0.245 0.758 34.2 0.220 0.828 54.6 0.389 0.731 48.5 0.318 0.680 43.5 0.290 0.803
PnP-ADMM 94.4 0.556 0.346 71.2 0.335 0.700 87.6 0.461 0.699 84.7 0.373 0.787
Score-SDE 104.2 0.573 0.543 48.1 0.304 0.649 105.7 0.596 0.451 149.8 0.625 0.294
ADMM-TV 151.3 0.458 0.692 72.3 0.287 0.812 133.2 0.502 0.671 118.6 0.492 0.704
C-DPS 28.6 0.178 0.779 29.9 0.205 0.835 50.1 0.345 0.735 45.6 0.300 0.688 41.2 0.263 0.823
C-DPS vs. Best -1.2 -0.067 0.012 -4.0 -0.015 0.007 -2.2 -0.022 -0.001 -1.7 -0.018 -0.021 -2.3 -0.027 0.006
Latent-Domain Methods
[…]
FFHQ PSLD 43.5 0.203 0.830 39.1 0.142 0.833 78.3 0.291 0.652 85.5 0.310 0.702 66.8 0.262 0.669
ReSample 36.9 0.126 0.776 49.1 0.166 0.779 59.4 0.234 0.736 41.1 0.189 0.845 83.6 0.344 0.619
LC-DPS 34.3 0.123 0.843 37.6 0.130 0.841 54.2 0.215 0.777 42.9 0.258 0.837 58.7 0.176 0.812
LC-DPS vs. Best -2.6 -0.003 0.013 -1.5 -0.012 0.008 -5.2 -0.019 0.041 1.8 0.069 -0.008 -8.1 -0.086 0.143
ImageNet PSLD 70.8 0.284 0.813 124.4 0.415 0.720 86.1 0.352 0.702 112.9 0.446 0.612 92.2 0.326 0.714
ReSample 49.1 0.118 0.782 111.9 0.223 0.659 57.3 0.225 0.724 61.5 0.200 0.745 107.2 0.332 0.609
LC-DPS 49.5 0.115 0.815 101.4 0.215 0.740 59.1 0.216 0.740 58.4 0.188 0.755 82.9 0.252 0.701
LC-DPS vs. Best 0.4 -0.003 0.033 -10.5 -0.008 0.020 1.8 -0.009 0.016 -3.1 -0.012 0.010 -9.3 -0.074 -0.013
H.2 More Qualitative Results
In this section, we present additional qualitative results on the FFHQ dataset across various inverse problem settings. These results serve to further demonstrate the flexibility and effectiveness of C-DPS in handling diverse measurement operators beyond those shown in the main text.
Figure˜6 shows reconstructions for the Gaussian deblurring task, where C-DPS successfully restores high-frequency details that are often lost in baseline methods. Figure˜7 presents results for motion deblurring, highlighting C-DPS’s ability to recover sharp facial features under structured blur.
In Figure˜8, we evaluate C-DPS on an 8× super-resolution task. Despite the aggressive downsampling, our method reconstructs globally consistent and perceptually plausible images. Figures 9 and 10 display results on random inpainting and box inpainting, respectively. In both cases, C-DPS generates semantically coherent completions even under extreme occlusions.
These qualitative results further support the claims made in the main paper, showing that C-DPS generalizes well across a wide range of inverse problems with minimal task-specific tuning.
Figure 6: Qualitative results on FFHQ dataset, Gaussian deblurring.
Figure 7: Qualitative results on FFHQ dataset, motion deblurring.
Figure 8: Qualitative results on FFHQ dataset, super-resolution task ($\times$ 8).
Figure 9: Qualitative results on FFHQ dataset, inpaiting task (random).
Figure 10: Qualitative results on FFHQ dataset, inpaiting task (box).
Appendix I Limitations and Failure Cases
While C-DPS demonstrates strong performance across a variety of inverse problems, we observe that it can struggle in certain edge cases. For example, when the measurement matrix $\mathbb{A}$ is nearly rank-deficient or the noise level is extremely high, the propagated measurement information $\{\boldsymbol{y}_{t}\}$ may not sufficiently constrain the posterior, resulting in oversmoothing or loss of fine details in the reconstruction.
This behavior is not unique to C-DPS. Prior methods such as DPS and ReSample also degrade under similar conditions, as they rely on learned priors and approximate measurement integration that become less reliable when the inverse problem is highly ill-posed. In our case, the Gaussian approximation to the posterior—while principled and efficient—can be limiting when the true posterior is strongly multi-modal or non-Gaussian.
These limitations arise from current design choices that favor tractability and generality. Future extensions could incorporate adaptive diffusion schedules, refined measurement dynamics, or non-Gaussian approximations to further improve performance in these extreme regimes.
Supplementary Materials (Evaluation on Medical Imaging Inverse Problems)
In this section, we extend our study by applying the proposed diffusion-based approach to inverse problems in the medical imaging domain. Medical images often contain complex structures and are typically acquired under constraints that make problems such as reconstruction, denoising, or inpainting particularly challenging. By evaluating our method on representative medical imaging tasks, we aim to demonstrate its effectiveness and generalization ability in real-world scenarios where accurate recovery of fine details is critical.
Datasets and Preprocessing.
We evaluate our proposed method, C-DPS, on three representative medical imaging tasks using public datasets: undersampled MRI, sparse-view CT, and super-resolution. For MRI reconstruction, we use the fastMRI knee dataset [zbontar2018fastmri]. Following standard preprocessing [song2022solving], we crop the raw k-space data to $320\times 320$ resolution and reconstruct single-coil images using a minimum variance unbiased estimator. Undersampling is simulated using 1D Gaussian and Uniform sampling masks with acceleration rates (ACR) of 4 and 8.
For sparse-view CT, we utilize the LIDC dataset [armato2011lung]. Two-dimensional CT slices of size $320\times 320$ are extracted from volumetric scans. Sinograms are generated using a parallel-beam geometry with either 23 or 10 projections evenly spaced over 180 degrees to simulate sparse-view acquisition.
The super-resolution task is evaluated on the fastMRI brain dataset [zbontar2018fastmri]. We downsample 2D full-resolution images to generate low-resolution inputs for $2\times 2$ and $4\times 4$ upscaling. We select approximately 63% of the slices labeled as reconstruction_rss, yielding a total of 34,698 training samples.
Baselines and Evaluation Metrics
We compare our method, C-DPS, with strong training-free diffusion-based baselines including DPS [DPS], Score-MRI [chung2022score], DDS [chungdecomposed], and ScoreMed [song2022solving], depending on the task. All methods are implemented with their publicly available code and evaluated under consistent experimental conditions.
We report results using standard image quality metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM). Each task is evaluated on a test set of 1,000 images. For CT experiments, we follow prior work and set the likelihood step size $\zeta=0$ where applicable. In all experiments, our method outperforms the baselines across all configurations in terms of both PSNR and SSIM, as shown in Tables 9, 11, and 10.
Architecture and Sampling Settings
All models are based on the ADM architecture [dhariwal2021diffusion], without classifier-free guidance or dropout. Separate diffusion priors are trained for each task. For sampling, we use 100 timesteps for MRI and super-resolution, and 350 timesteps for sparse-view CT.
Our method introduces a measurement-consistent update through a bi-level guidance scheme. We tune the likelihood step size $\zeta$ and refinement weight $\gamma$ within a restricted search space, $\zeta\in[0,2]$, $\gamma\in[0,4.5]$, and fix $\lambda=10^{-3}$ for the outer optimization. No extensive hyperparameter tuning is required to achieve strong performance.
All experiments are conducted on a NVIDIA P100 GPU with 12 GB of memory.
Results
We assess the effectiveness of our method, C-DPS, across three key inverse problems: MRI reconstruction, CT reconstruction, and super-resolution. In all cases, C-DPS consistently outperforms prior training-free diffusion-based baselines, demonstrating both robustness and accuracy across varying acquisition settings.
MRI Reconstruction.
[…]
Sparse-View CT Reconstruction.
As shown in Table 11, C-DPS achieves state-of-the-art results on the LIDC dataset with both 23 and 10 projections. It outperforms ScoreMed and BGDM by a notable margin in PSNR and SSIM, despite using the same number of diffusion steps (350). The improvement is especially pronounced under the extremely sparse 10-view setting, suggesting that C-DPS maintains strong data consistency even under severe measurement limitations.
Super-Resolution.
On the fastMRI brain dataset, Table 10 demonstrates that C-DPS achieves the highest reconstruction quality for both $\times2$ and $\times4$ super-resolution tasks. Compared to DPS and DDS, our method produces sharper images with higher structural similarity, particularly in the more difficult $\times4$ setting. These results affirm the generalization capability of C-DPS across different types of inverse problems.
Overall Observations.
Across all tasks, C-DPS improves upon existing baselines without requiring task-specific training or extensive parameter tuning. The results highlight the advantage of incorporating measurement-aware updates through our bi-level guidance framework, which enhances both reconstruction accuracy and robustness under limited data.
Table 9: Quantitative results for the fastMRI knee dataset across different sampling masks and acceleration rates.
[…]
$uparrow PSNR\uparrow SSIM\uparrow$ $uparrow PSNR\uparrow SSIM\uparrow$
[…]
Table 10: Super-resolution results on the fastMRI brain dataset.
Method $\times2$ SR $\times4$ SR
$uparrow PSNR\uparrow SSIM\uparrow$
[…]
Table 11: Quantitative results of sparse-view CT reconstruction on the LIDC dataset using 350 NFEs.
Method 23 Projections 10 Projections
$uparrow PSNR\uparrow SSIM\uparrow$
[…]
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We provide a clear summary of our results in the abstract and introduction, stating all conditions needed for the results to hold. We summarize the experiments presented in the paper, with details in the main section and the Appendix.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discusses all the necessary conditions for the results to hold. In the last section of the paper, we discuss multiple limitations of the proposed approach.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The justification/proof for the theoretical parts are provided in the main body and appendix of the paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We describe details of the model, including hyperparameters and system prompts, in the paper and in the appendix. Moreover, we provide the link to the anonymous repository that contains all of the codes needed to run the experiments demonstrated in the paper and it contains the link to the dataset proposed in the paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the link to the anonymous repository that contains all of the codes needed to run the experiments demonstrated in the paper
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes] .
  29. Justification: We provide the experimental setting in the paper and also more details in the appendix. We present all the necessary to run the experiments together with the code on the anonymous repository.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: While our main results are based on extensive evaluation across multiple datasets and inverse problem settings, we do not report error bars or confidence intervals in the main figures and tables. However, our method consistently outperforms baselines across all tested configurations, and we have verified the stability of our results over multiple random seeds.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the details of the machine where we performed experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The paper conforms to the NeurIPS Code of Ethics. In particular, the datasets used are open widely used in related work; we do not use human participants
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our paper focuses on a theoretical and algorithmic advancement in solving inverse problems using diffusion models. While we do not explicitly discuss societal impacts, the method has potential positive applications in areas such as medical imaging, where improved reconstructions can aid diagnosis. At the same time, as with any generative model, there is a possibility of misuse in domains like surveillance or data manipulation. We acknowledge the importance of broader impact assessments and encourage responsible deployment of such methods.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite all the datasets used, noting that they are open for use. We cite any code used for our implementations.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We describe all datasets used, with proper documentation of parameters used, algorithms, and code instructions.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not involve crowdsourcing or research with human subjects
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs are used only for writing purposes.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.